### v1

Abstract
The scaling law, a cornerstone of Large Language Model (LLM) development, predicts improvements in model performance with increasing computational resources. Yet, while empirically validated, its theoretical underpinnings remain poorly understood. This work formalizes the learning dynamics of transformer-based language models as an ordinary differential equation (ODE) system, then approximates this process to kernel behaviors. Departing from prior toy-model analyses, we rigorously analyze stochastic gradient descent (SGD) training for multi-layer transformers on sequence-to-sequence data with arbitrary data distribution, closely mirroring real-world conditions. Our analysis characterizes the convergence of generalization error to the irreducible risk as computational resources scale with data, especially during the optimization process.
We establish a theoretical upper bound on excess risk characterized by a distinct phase transition. In the initial optimization phase, the excess risk decays exponentially relative to the computational cost $\sf C$}. However, once a specific resource allocation threshold is crossed, the system enters a statistical phase, where the generalization error follows a power-law decay of $\Theta(\mathsf{C}^{-1/6})$. Beyond this unified framework, our theory derives isolated scaling laws for model size, training time, and dataset size, elucidating how each variable independently governs the upper bounds of generalization.
1 Introduction
The emergence of transformer-based Large Language Models (LLMs) such as ChatGPT [27], GPT-4 [1, 18], DeepSeek [17, 50, 49, 33], and LLaMA [71, 72, 33] has driven transformative advancements across multiple domains [22]. Tasks like code generalization [47, 35], conversational systems [56, 78, 84], and mathematical reasoning [37, 82, 83], once considered exclusive to human expertise, are now routinely mastered by these AI systems.
This remarkable progress is fundamentally tied to computational scaling. Empirical evidence reveals a consistent pattern: as the compute budget for optimally training and deploying language model increases, their demonstrated intelligence scales correspondingly [44, 38]. This phenomenon has been systematically categorized into the following principle:
Scaling Law [44, 38]: Model capabilities improve predictably with increased computational investment during training, achieved through three-dimensional scaling: model parameter count, training duration, and dataset size.
While the scaling law has been extensively validated through empirical research, a critical gap persists in its theoretical foundation. Current literature lacks a mathematically rigorous framework to explain why these principles exhibit such predictable improvements in model performance, or to systematically justify their reliability in guiding the development of massive-scale transformer architectures [54]. Existing theoretical studies on scaling laws have predominantly focused on traditional statistical models and simplified neural architectures. Key investigations include analyses of shallow attention networks [55], linear regression [54, 30], kernel regression [26], and data selection methodologies [43], among others. However, these works lack generalizability to modern deep learning systems. Notably, the scaling theory underpinning transformer-based LLMs [74] — the dominant paradigm in contemporary AI — remains largely unexplored.
Our work addresses this gap by establishing theoretical foundations for computational scaling effectiveness in training multi-layer transformer-based language models on sequence-to-sequence data. Specifically, we derive a quantifiable relationship between computational expenditure and model capabilities, formulating an optimized error bound for training objectives that explicitly depends on allocated computational resources. In particular, the goal of this work is to address the following critical question:
What are the theoretical limits for computational resource allocation to ensure the convergence of the generalization error bound during transformer-based language model training?
We present the first comprehensive analysis of the training dynamics of transformer-based language models using the Stochastic Gradient Descent (SGD) algorithm, along with a convergence guarantee for arbitrary training error. This topic has been largely unexplored due to the inherent complexity of attention mechanisms, the multi-layered structure of transformers, and the extensive matrix computations required for sequence-to-sequence data processing. To address these challenges, we adopt a kernel-based analytical framework to investigate the scaling behavior. By analyzing this scenario, we establish the lower and upper bounds for the expected risk on the whole data distribution. Overall, we make the following contributions:
  • We first simplify the complicated matrix computation to the parallel vector computation utilizing the decoder-only property of a generative transformer-based language model. Therefore, we formalize its explicit learning dynamics, which we assume the learning behavior of LLMs is constrained to the kernel regimes due to the over-parameterization. This is referred to the Lazy Learning [42, 31], where the model merely memorizes all data points during optimization. (See Section 4 and Section 6)
  • We further explore several benefits that emerge under the trend of model scaling, e.g., converging kernel perturbation, which means that once kernel behavior exists at initialization, it will remain stable during training. Moreover, we demonstrate how the training convergence rate exponentially improves with linearly increasing model depth. Finally, we showcase the guaranteed training convergence and generalization bound of the empirical risk minimizer (ERM) related to vital scaling factors (model size, dataset size). (See Section 5)
  • The primary contribution of this work is a two-stage upper bound on the in-distribution expected risk, which formalizes a phase transition between a compute-starved stage and a data-limited stage. Our analysis reveals that the generalization error initially undergoes an exponential decay governed by total computational expenditure, $\sf C$}, during which the expected risk is dominated by optimization dynamics. Upon reaching a critical saturation point, defined by the condition in Eq. (3), the scaling trajectory undergoes a structural shift: the rapid exponential decrease transitions into a power-law regime. In this subsequent stage, the frontier error is fundamentally constrained by the target function’s intrinsic complexity and the dataset size, $\sf N$}. Consequently, further increases in training duration or model capacity yield diminishing returns on generalization performance unless accompanied by a proportional expansion of high-quality data. (See Section 7)
2 Related Work
Scaling Law. Several recent works have empirically explored scaling laws in deep neural networks [44, 38, 60, 41, 61]. The study of neural scaling laws can also be traced back to earlier foundational works [25, 64, 8]. From a theoretical perspective, various solvable models have been developed using random feature models [20, 2, 15, 16, 59] to analyze neural scaling laws under specific limits. Additionally, theoretical analyses on linear models [75, 19, 65, 23, 54, 55] have significantly advanced our understanding of scaling laws. In contrast to these studies, our work focuses on the sequence-to-sequence stochastic training of multi-layer transformer-based language models, a topic that has not been widely discussed in prior research.
Neural Tangent Kernel and Learning Theory. The Neural Tangent Kernel (NTK), introduced by [42], has become a foundational framework for understanding the gradient flow of neural networks during training. It reveals that, in the infinite-width limit, neural networks are equivalent to Gaussian processes at initialization. This equivalence has been extensively studied in numerous works [51, 31, 70, 14, 76, 21, 53, 24, 68, 85, 65, 34, 52, 69], which highlight the robust performance and learning capabilities of over-parameterized neural networks. The NTK framework has gained popularity for its ability to elucidate the emerging abilities of large-scale neural networks. Notable advancements include the introduction of the Convolutional NTK (CNTK) by [3], the Recurrent NTK by [9], and the concept of infinite attention via NNGP and NTK for attention networks by [39]. Furthermore, [57] examined the training dynamics of fine-tuning LLMs using NTK, demonstrating its efficiency in optimizing these systems. These contributions underscore the growing importance of NTK in the theoretical analysis of modern neural networks.
Science of Transformer-based Language Models. The complex architecture and stochastic optimization processes of transformer-based language models pose significant challenges for theoretical analysis. However, developing theoretical guarantees for LLMs is essential for advancing their design and performance. Recent research has addressed various aspects of LLMs, including improving efficiency [5, 6, 7, 40, 45, 4, 29, 66], optimizing training processes [28, 52], analyzing ”white-box” transformers [80, 81, 32, 58], and investigating the emergent abilities of LLMs [22, 77, 10, 12, 11, 13]. By bridging theoretical understanding with practical advancements, these studies provide valuable insights for the development of the next generation of AI systems.
3 Preliminary
This section provides the preliminary for our analysis, where we introduce some basic notations in Section 3.1, and present the problem setup in Section 3.2. We encourage the reader to refer to Appendix A for the formal technical preliminary.
3.1 Basic Notations
Let $d]=\{1,2,\dots,d\}$. For $u\in\mathbb{R}^{d}$, the $\ell_{p}$-norm is $\|u\|_{p}:=(\sum_{k=1}^{d}|u_{k}|^{p})^{1/p}$. The Frobenius norm of $U\in\mathbb{R}^{d_{1}\times d_{2}}$ is $\|U\|_{F}:=(\sum_{(k_{1},k_{2})\in[d_{1}]\times[d_{2}]}U_{k_{1},k_{2}}^{2})^{1/2}$. For a matrix $A\in\mathbb{R}^{d\times d}$, $\lambda_{\min}(A)$ denotes its smallest eigenvalue. The indicator function $\operatorname{\mathbb{I}}\{E_{1},\dots,E_{n}\}$ equals $1$ if all events $E_{1},\dots,E_{n}$ occur, and 0 otherwise. The mapping $\operatorname{mat}:\mathbb{R}^{d^{2}}\to\mathbb{R}^{d\times d}$ reshapes a vector $a$ into a matrix such that $\operatorname{mat}_{k_{1},k_{2}}(a)=a_{(k_{1}-1)d+k_{2}}$ for $(k_{1},k_{2})\in[d]\times[d]$. Conversely, $\operatorname{vec}:\mathbb{R}^{n\times d}\to\mathbb{R}^{nd}$ flattens a matrix $A$ into a vector with $\operatorname{vec}_{k}(A)=A_{\lfloor k/d\rfloor,\,k-\lfloor k/d\rfloor\cdot d}$ for $k\in[nd]$. For a function $f:X\to\mathbb{R}^{d_{1}\times d_{2}}$, $f_{k_{1},k_{2}}(x)$ denotes the $(k_{1},k_{2})$-entry of $f(x)$. $e_{k}\in\mathbb{R}^{d}$ is the standard basis vector with a $1$ in the $k$-th entry and 0 elsewhere. Finally, we denote $a\wedge b:=\max\{a,b\}$ for $a,b\in\mathbb{R}$.
3.2 Setups
[…]
$~\text{Expected Risk: }{\cal R}(F):=\operatorname*{{\mathbb{E}}}_{(X,Y)\sim{\cal D}}[\|F(X)-Y\|_{F}^{2}],\quad\text{Excess Risk: }\Delta{\cal R}(F):={\cal R}(F)-{\cal R}(F^{*}).$
Besides, we have an accessable dataset $\mathbb{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}\subset{\cal D}$ where each data point independently and uniformly sampled from $\cal D$}. The random noise $\Xi\in\mathbb{R}^{L\times d}$ is centered by ${\bf 0}_{L\times d}$ with variance $\max_{\ell\in[L]}\max_{k\in[d]}\operatorname*{\mathrm{Var}}[\Xi_{\ell,k}]=\xi^{2}$. For any input matrix $X\sim{\cal X}$, $\|X_{\ell}\|_{2}=\Theta(1)$ holds for each token vector $\ell\in[L]$ due to the utilization of RMS normalization [86] after the embedding layer.
Model Function. The standard transformer architecture introduced in [74] integrates multiple self-attention layers with token-wise feed-forward layers. The fundamental architecture, decoder-only transformers [62], processing a sequence of $L$ tokens, each represented by a $d$-dimensional embedding vector, which are compactly arranged into a matrix $X\in\mathbb{R}^{L\times d}$. An $N$-layer transformer model is formally defined as:
$F(X,\theta):=\varepsilon\cdot F_{(N)}(F_{(N-1)}(\cdots F_{(2)}(F_{(1)}(X+E,\theta),\theta)\cdots),\theta),$ (1)
where $E\in\mathbb{R}^{L\times d}$ is the positional embedding matrix [^3] We choose $E={\bf 0}_{L\times d}$ (NoPE, No Positional Embedding, [46]) or ignore it as a fixed matrix (this can be regarded as a part of the training dataset) in the range of this paper. and $\theta$ is the set of all trainable parameters. $\varepsilon>0$ is the scaling coefficient to control the norm of the initial parameters, which helps to reduce the model complexity in the analysis. Each $F_{(\nu)}$ (for $\nu\in[N]$) represents the $\nu$-th transformer block and is given by:
$F_{(\nu)}(X,\theta):=\frac{\omega}{\sqrt{m}}{\rm ReLU}\left({\rm Softmax}\left(\kappa\cdot XU_{(\nu)}X^{\top}+M\right)XW_{(\nu)}\right)A_{(\nu)}+X,$
where $U_{(\nu)}\in\mathbb{R}^{d\times d}$, $W_{(\nu)},A_{(\nu)}^{\top}\in\mathbb{R}^{d\times m}$ are model parameters. $\kappa$ is the scaling factor of attention, $\omega$ is the scaling coefficient of output. $M\in\mathbb{R}^{L\times L}$ is the causal attention mask. We especially use $w_{(\nu),r},a_{(\nu),r}\in\mathbb{R}^{d}$ to denote the $r$-th column of $W_{(\nu)}$ and the $r$-th row of $A_{(\nu)}$, respectively.
Initialization and Training. For every layer $\nu\in[N]$, the corresponding parameters of $F_{(\nu)}$ is denoted as $U_{(\nu)},W_{(\nu)},A_{(\nu)}$. Each entry of $U_{(\nu)},W_{(\nu)}$ is initialized from the standard Gaussian distribution $\mathcal{N}(0,1)$, we then denote them as $U_{(\nu)}(0),W_{(\nu)}(0)$. Besides, each entry of $A_{(\nu)}$ is initialized from a uniform distribution ${\sf Uniform}\{-1,+1\}$ and is frozen during training. The flattened vector of the whole trainable parameters is denoted as $\theta(0)\in\mathbb{R}^{{\sf M}}$ where ${\sf M}=N(md+d^{2})$ is the number of trainable parameters.
Given the training dataset $\mathbb{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}\subset{\cal D}$, we define the overall training objective as:
${\cal L}(t,\mathbb{D}):=\operatorname*{{\mathbb{E}}}_{(X,Y)\sim\mathbb{D}}[\|F(X,\theta(t))-Y\|_{F}^{2}].$
Therefore, we consider a combination of the stochastic gradient descent (SGD) algorithm and gradient flow to update. At $t$-step optimization, we sample a unbiased subset $\mathbb{B}(t)\subseteq\mathbb{D}$ satisfying $\operatorname*{{\mathbb{E}}}[{\cal L}(t,\mathbb{B}(t))]={\cal L}(t,\mathbb{D})$ for training time $T>0$. We let the batch size $|\mathbb{B}|={\alpha}_{\rm agr}\cdot n$ fixed with some proportion ${\alpha}_{\rm agr}\in(0,1]$, its value is arbitrary to set due to the trick of accumulative gradient. Then the ordinary differential equation (ODE) of $U_{(\nu)}(t),W_{(\nu)}(t)$ and their update rule are given by:
$\begin{split}&~\frac{\mathrm{d}}{\mathrm{d}t}U_{(\nu)}(t)=-\frac{\mathrm{d}}{\mathrm{d}U_{(\nu)}(t)}{\cal L}(t,\mathbb{B}(t)),\quad\frac{\mathrm{d}}{\mathrm{d}t}W_{(\nu)}(t)=-\frac{\mathrm{d}}{\mathrm{d}W_{(\nu)}(t)}{\cal L}(t,\mathbb{B}(t)),\\ &~U_{(\nu)}(t+\tau)=U_{(\nu)}(t)+\int_{t}^{t+\tau}\frac{\mathrm{d}}{\mathrm{d}s}U_{(\nu)}(s)\mathrm{d}s,\quad W_{(\nu)}(t+\tau)=W_{(\nu)}(t)+\int_{t}^{t+\tau}\frac{\mathrm{d}}{\mathrm{d}s}W_{(\nu)}(s)\mathrm{d}s.\end{split}$ (2)
Hence, we denote the training algorithm that depends on training time and dataset size as ${\cal A}_{T,n}(\theta(0),$ $\mathbb{D})$ $\{\theta(0)+\int_{0}^{T}-\frac{\mathrm{d}}{\mathrm{d}\theta}{\cal L}(s,\mathbb{B}(s))\mathrm{d}s,\mathbb{B}(s)\subseteq\mathbb{D},s\in[0,T]\}$.
4 Learning Dynamics of Scaling Transformers
In this section, we provide the explicit model learning dynamics formulations layer by layer. We first introduce several key simplifications in Section 4.1 to facilitate the preliminary analysis. Thus, Section 4.2 formulates an explicit ODE of the learning dynamics of scaling multi-layer transformer-based language models.
4.1 Simplifications
Utilizing the attention network’s decoder-only property, it is obvious that the matrix computation can be parallelized to vector computation. We use $X_{i,\leq\ell}\in\mathbb{R}^{\ell\times d}$ to denote the first-$\ell$ tokens ($\forall\ell\in[L]$) of matrix $X_{i}$ in $\mathbb{D}$. Hence, we compact the outputs of the model function and targets on the whole dataset to two matrices ${\sf F}(t),{\sf Y}\in\mathbb{R}^{nL\times d}$, where ${\sf F}_{(i-1)L+\ell}(t)=F_{\ell}(X_{i},\theta(t))\in\mathbb{R}^{d}$ and ${\sf Y}_{(i-1)L+\ell}=Y_{i,\ell}$ for each $(X_{i},Y_{i})$ in training dataset $\mathbb{D}$ and $\ell\in[L]$. Therefore, we derive ${\cal L}(t,\mathbb{D})=\frac{1}{n}\|{\sf F}(t)-{\sf Y}\|_{F}^{2}$ (See Lemma B.3). Besides, we list following helpful functions ($(i,\ell)\in[n]\times[L]$):
  • (Hidden State) $\Lambda_{(\nu),i}(t):=F_{(\nu)}(\Lambda_{(\nu-1),i}(t),\theta(t))\in\mathbb{R}^{L\times d}$ for $\nu\in[N]$, $\Lambda_{(0),i}(t)=X_{i}+E$.
  • (Attention Scores) $\sigma_{(\nu),(i-1)L+\ell}(X)={\rm Softmax}_{\ell}(\Lambda_{(\nu),i}(t)U_{(\nu)}(t)\Lambda_{(\nu),i}(t)^{\top}+M)\in\mathbb{R}^{L}$.
  • (Attention Output) $o_{(\nu),(i-1)L+\ell}(t):=\Lambda_{(\nu-1),i}(t)^{\top}\cdot\sigma_{(\nu),(i-1)L+\ell}(t)\in\mathbb{R}^{d}$.
  • ($\ell$-th Token of Hidden State)
    $\mu_{(\nu),(i-1)L+\ell}(t):=\frac{\omega}{\sqrt{m}}\sum_{r=1}^{m}a_{(\nu),r}\cdot\phi(\langle o_{(\nu),(i-1)L+\ell}(t),w_{(\nu),r}(t)\rangle)\in\mathbb{R}^{d},$
    where $\phi(x):=\max\{0,x\},\forall x\in\mathbb{R}$. $\mu_{(0),(i-1)L+\ell}(t)=X_{i,\ell}+E_{\ell}$.
  • (Model Output) ${\sf F}_{(i-1)L+\ell}(t)=\varepsilon\cdot\sum_{\nu=0}^{N}\mu_{(\nu),(i-1)L+\ell}(t)\in\mathbb{R}^{d}$.
4.2 Key Derivation for Learning Dynamics
The primary challenge in understanding the learning dynamics of finite-deep transformers is the complex analysis of gradient flow, which differs from the study of shallow or infinite-deep neural networks. We overcome the complexity by cleverly utilizing the multivariable chain rules.
First, we define the kernel matrix at $\nu$-th layer as $H_{(\nu)}\in\mathbb{R}^{nL\times nL}$ and its $(i,j)$-th entry ($\forall(p,q)\in[nL]\times[nL]$) is defined as:
$H_{(\nu),p,q}(t):=\underbrace{\langle\beta_{(\nu),p}(t),\beta_{(\nu),q}(t)\rangle}_{\text{kernel w.r.t. $W_{(\nu)}(t)$}}+\underbrace{\langle\gamma_{(\nu),p}(t),\gamma_{(\nu),q}(t)\rangle}_{\text{kernel w.r.t. $U_{(\nu)}(t)$}},$
Here, we let:
$\beta_{(\nu),p}(t):=$ $~\frac{\omega}{\sqrt{m}}\underbrace{o_{(\nu),p}(t)}_{d\times 1}\otimes\underbrace{{\bf 1}_{W_{(\nu)}(t)^{\top}o_{(\nu),p}(t)>0}}_{m\times 1}\in\mathbb{R}^{md},$
$\gamma_{(\nu),p}(t):=$ $~\frac{\omega\cdot\kappa}{\sqrt{m}}\underbrace{(\Lambda_{(\nu-1),i,\ell,*}(t)\otimes\Lambda_{(\nu-1),i}(t))^{\top}}_{d^{2}\times L}\underbrace{\left(\operatorname{diag}(\sigma_{(\nu),p}(t))-\sigma_{(\nu),p}(t)\sigma_{(\nu),p}(t)^{\top}\right)}_{L\times L}$
$~\underbrace{\Lambda_{(\nu-1),i}(t)}_{L\times d}\sum_{r\in[m]}\underbrace{w_{(\nu),r}(t)\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0\}}_{d\times 1}\in\mathbb{R}^{d^{2}},$
where $\otimes$ is the Kronecker product and $i=\lfloor p/L\rfloor$, $\ell=p{~\rm mod~}L$. The indicator vector ${\bf 1}_{W_{(\nu)}(t)^{\top}o_{(\nu),p}(t)>0}$ $\in\{0,1\}^{m}$ where its $r$-th entry is $\operatorname{\mathbb{I}}\{\left(W_{(\nu)}(t)^{\top}o_{(\nu),p}(t)\right)_{r}>0\}$ for $r\in[m]$. The layer-wise training dynamics are thereby shown as the following lemma:
Lemma 4.1 (Learning dynamics, informal version of Lemma B.6).
The learning dynamics of the multi-layer transformer Eq. (1) is given by:
$\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]=-\sum_{\nu\in[N]}{\underbrace{\operatorname{vec}\left(\frac{\mathrm{d}}{\mathrm{d}\mu_{(\nu)}(t)}{\cal L}(t,\mathbb{D})\right)^{\top}}_{1\times nLd}}\cdot\underbrace{\left(H_{(\nu)}(t)\otimes I_{d}\right)}_{nLd\times nLd}\cdot\underbrace{\operatorname{vec}\left(\frac{\mathrm{d}}{\mathrm{d}\mu_{(\nu)}(t)}{\cal L}(t,\mathbb{D})\right)}_{nLd\times 1}$
where $\mu_{(\nu)}(t)$ is a $nL\times d$ matrix, $\mu_{(\nu),p}(t)$ is the $(p\text{ mod }L)$-th row of $v$-th layer output regarding to input matrix $X_{\lfloor p/L\rfloor}$ for any $p\in[nL]$ and $\nu\in[N]$.
Proof sketch of Lemma 4.1.
Although the derivation of the gradient is complicated, the technique used for the proof is just the chain rule. We provide the complete proof in Appendix B. ∎
Lemma 4.1 sums contributions from all layers $\nu\in[N]$, providing a powerful decomposition. This granular view suggests that each layer independently contributes to minimizing the loss based on its own kernel and local gradient. Different layers may learn at varying speeds or contribute differently to the overall task, depending on their respective kernel structures.
5 Convergence of Transformers Training
In this section, we showcase the training convergence with an arbitrary error by limiting the kernel matrix $H_{(\nu)}^{\prime}(0)$ at initialization to the Neural Tangent Kernel (NTK) [42, 31] assumption. We state the formal assumption and the basic inductive proof in Section 5.1, and Section 5.2 demonstrates the results of the training convergence.
5.1 Assumptions and Inductions
Following the setting of [31] (see Assumption 3.1), we align this mild assumption in our analysis as shown:
Assumption 5.1.
Defining $H_{(\nu)}^{\prime}(0)\in\mathbb{R}^{nL\times nL}$ where its $(i,j)$-th entry ($\forall(i,j)\in[nL]\times[nL]$) is given by $H_{(\nu),i,j}^{\prime}(0):=\langle\beta_{(\nu),i}(t),\beta_{(\nu),j}(t)\rangle$. For all $\nu\in[N]$, we assume $\frac{1}{\omega}H_{(\nu)}^{\prime}(0)$ is positive definite (PD), formally, $\lambda_{(\nu)}:=\lambda_{\min}(\frac{1}{\omega}H_{(\nu)}^{\prime}(0))>0$.
Converging Kernel Perturbation and Kernel-Based Lazy Learning. Assuming $\lambda_{\min}(H_{(\nu)}^{\prime}(0))>0$, we extend this PD property to the case during the training and the case of $H_{(\nu)}(t)$ that requires bounded summational update of weight at an arbitrary positive time $t>0$. This connects to the Lazy Learning regime [42, 31], where the updates of weights are limited in some high-dimensional ball with radius $R$. We first give the Good Properties requirements for provable arbitrary convergence below:
Definition 5.2 (Good Properties and Good Model Class).
We fix the dataset size $n$, then we say model $F(X,\theta)$ has Good Properties once it satisfies (constant $C>0$ is some considerably small constant):
  1. 1.
    $\omega=\Theta(\frac{1}{NL^{2}d^{2.5}B^{3}})$;
  2. 2.
    $\kappa=\frac{1}{\sqrt{m}}$;
  3. 3.
    $m=\Omega(\frac{n^{3}L^{5}\exp(Cd)}{\omega^{6}\lambda^{6}\delta^{3}N^{2}})$.
Good Model Class is ${\cal F}_{\sf M,T,N}(\mathbb{D}):=\{F(\cdot,\theta({\sf T})),\theta(0)\sim\mathcal{N}(0,I_{\sf M}),\theta({\sf T})\in{\cal A}_{{\sf T,N}}(\theta(0),\mathbb{D})\}$ for any $\mathbb{D}\subset{\cal D}$ and $|\mathbb{D}|={\sf N}$.
Lemma 5.3 (Informal version of Lemma D.1).
Assuming Assumption 5.1 and Definition 5.2 hold, denote the failure probability $\delta\in(0,0.1)$, then the kernel perturbation bound is:
$\Pr\left[\lambda_{\min}(H_{(\nu)}(t))<\lambda/2\right]<\delta.$
Therefore, bounding loss dynamics is given by ($C>0$ is some constant):
$\Pr\left[\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]>-C\cdot\omega\lambda N\cdot{\cal L}(t,\mathbb{D})\right]<\delta.$
Proof sketch of Lemma 5.3.
The stability of the PD property is ensured by $\lambda_{\min}(H^{\prime})\geq\lambda_{\min}(H)-\|H-H^{\prime}\|_{F}$ (Fact A.11) since the considerably small perturbation $\|H-H^{\prime}\|_{F}$. The upper bound on perturbation converges due to increasing feed-forward layer width $m$ and decreasing weight perturbation radius $R$, leading to the Lazy Learning regime. The complete proof of this lemma is stated in Appendix D. ∎
5.2 Convergence under Kernel Regimes
Factors of Scaling Law. To begin with, we list several crucial factors of the scaling law below. We ignore the batch size as a constant in the analysis of the total compute. In addition, the model size refers to the number of trainable parameters in our model, which could be trivially calculated, and the definition of the total compute holds as our compute analysis in Lemma E.1.
Definition 5.4.
We define:
  • Model Size: ${\sf M}:=O(N(md+d^{2}))=O(Nmd)$.
  • Dataset Size: ${\sf N}:=n$.
  • Training Time: ${\sf T}:=T$.
  • Total Compute: ${\sf C}:=O({\sf MT}|\mathbb{B}|)\leq O({\sf MTN})$.
Training Convergence. The convergence guarantee of implementing Eq (2) as the continuous-time optimization is stated below:
Theorem 5.5 (Informal version of Theorem E.2).
Let all scaling law factors be defined as Definition 5.4 and Assumption 5.1 and Definition 5.2 hold. Denote the failure probability $\delta\in(0,0.1)$, $\alpha_{\rm cr}=O(\lambda/\omega)$ (maximum convergence rate). For the Good Model Class ${\cal F}_{\sf M,T,N}(\mathbb{D})$, with a probability at least $1-\delta$, we have:
${\cal L}({\sf T},\mathbb{D})\leq\exp(-\alpha_{\rm cr}\cdot\varepsilon^{2}T)\cdot{\cal L}(0,\mathbb{D}),\forall F\in{\cal F}_{\sf M,T,N}(\mathbb{D}).$
Proof Sketch of Theorem 5.5.
Firstly, we confirm the connection between hidden-state gradient norm and the training objective in Part 15 of Lemma C.1, the model convergence therefore exponentially benefits from the model size (neural depth and width) and the training time. Besides, the variance produced by the stochastic algorithm is provably reduced with a considerably large $m$. See Appendix E.2 for the detailed proofs. ∎
6 Generalization Bound of Empirical Risk Minimizer
This section formally analyzes the generalization error bound of transformer-based language models. In particular, Section 6.1 provides the approximation error w.r.t the model size $\sf M$} via the approximation capability of kernel regression in RKHS (Reproducing Kernel Hilbert Space), and Section 6.2 showcases the general scaling law, which describes how the generalization error bound converges with the growing total computational cost.
6.1 Approximation
In order to estimate the upper bound on the excess risk, we break it down into approximation error, generalization (estimation) error, and irreducible error [67, 48]. We first give the bound on approximation below, which measures the model’s best capability to approximate the target function on the entire input space.
Corollary 6.1 (Informal version of Corollary E.3).
Assuming we have arbitrary dataset size ${\sf N}\in(0,+\infty)$. Let all scaling law factors be defined as Definition 5.4 and Assumption 5.1 and Definition 5.2 hold. Denote the failure probability $\delta\in(0,0.1)$. For the Good Model Class ${\cal F}_{\sf M,T,N}(\mathbb{D})$ with training dataset $\mathbb{D}$ and target function $F^{*}:{\cal X}\to{\cal Y}$, we choose $\varepsilon\lesssim{\sf M}^{-1}$, with a probability at least $1-\delta$, we have:
$~\lim_{{\sf T}\rightarrow+\infty}\inf_{F\in{\cal F}_{\sf M,T,N}(\mathbb{D})}\operatorname*{{\mathbb{E}}}_{(X,Y)\sim\mathcal{D}}[\|F(X)-F^{*}(X)\|_{F}^{2}]\leq{\sf M}^{-2}.$
Proof Sketch of Corollary 6.1.
According to the conditions of the target function $F^{*}$ in Section 3, we formulate its expression in the RKHS (Reproducing Kernel Hilbert Space), hence, the approximation error will be bounded by a considerably small $\varepsilon$ and over-parameterized model width $m$, which is similar to Theorem 3.2 in [3]. The formal proof is in Appendix E.3. ∎
6.2 Generalization
Here, we answer the question of how well a transformer-based empirical risk minimizer (ERM) performs statistical generalization, which, in detail, we scale the training time $\sf T$} till infinity to get an upper bound on excess risk $\Delta{\cal R}(F)$.
Theorem 6.2 (Informal version of Theorem F.6).
Let all pre-conditions hold as Corollary 6.1 and choose $\varepsilon=\Theta(1/{\sf M})$. We denote $\inf_{{\cal F}_{\sf M,N}(\mathbb{D})}:=\lim_{{\sf T}\to\infty}\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}$ Then, with a probability at least $1-\delta$, there exists:
$\inf_{{\cal F}_{\sf M,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq(4{\sf M}^{-1}+\xi)\cdot{\sf M}^{-1}+\frac{Ld\cdot\xi}{\sf N}.$
Proof Sketch of Theorem 6.2.
This proof utilizes the results in Lemma 11 of [63] for the generalization error and Corollary 6.1 for the approximation error. The full proof is shown in Appendix F.1. ∎
7 Scaling Law and Its Potential Drag
In this section, we unify our results from Section 5 (training convergence dynamics) and Section 6 (generalization bound of ERM) into one. We derive our theory of transformer scaling law with the discovery of a pattern transition inside in Section 7.1. Next, in Section 7.2, we discuss several vital factors that can considerably affect the success of the scaling process. For every situation that might lead to the failure of scaling, we call it potential drag of the scaling law.
7.1 General Scaling Law with Pattern Transition
Theorem 7.1 (Informal version of Theorem F.7).
Let all pre-conditions hold as Corollary 6.1 and choose $\varepsilon=\Theta(\xi/{\sf N})$. Then, with a probability at least $1-\delta$, the following condition divides the scaling trend into two stages:
${\sf C}>\frac{{\sf N}^{6}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$ (3)
  • Stage I: Compute-Starved Stage. If Eq. (3) does not hold for any dataset size ${\sf N}>0$, we have:
    $C\exp(-\alpha\xi^{2}{\sf C}/{\sf N}^{6})\cdot{\cal L}(0,\mathbb{D}),$
    where the constant $C=\operatorname{poly}(L,d,1/\lambda,1/\omega)$.
  • Stage II: Data-Limited Stage. If Eq. (3) holds for any dataset size ${\sf N}>0$, we have:
    $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq\xi^{5/3}\cdot O\left(\frac{{\sf C}}{W({\sf C}/\xi^{10})}\right)^{-1/6}\leq\xi^{5/3}\cdot\widetilde{O}\left({\sf C}\right)^{-1/6},$
    where $W(\cdot)$ is the Lambert W function and $\widetilde{O}$ hides $\log({\sf C})$.
Proof Sketch of Theorem 7.1.
We decompose the excess risk $\Delta{\cal R}(F)$ into three terms: i) optimization risk ${\cal R}_{\rm optimization}:=\|F(\cdot,\theta(t))-F(\cdot,\theta(\infty))\|_{L^{F}({\cal X})}^{2}$; ii) approximation risk ${\cal R}_{\rm approximation}:=\left\|F(\cdot,\theta(\infty))-{\rm arg}\inf_{F^{\prime}\in{\cal F}_{\sf M,\infty,N}}\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}^{2}\right\|_{L^{F}({\cal X})}^{2}$; and iii) estimation risk ${\cal R}_{\rm estimation}$ $\inf_{F^{\prime}\in{\cal F}_{\sf M,\infty,N}}$ $\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}^{2}$. The second and the third terms are already bounded in Theorem 6.2, and the first term relates to the training loss; we thereby utilize Theorem 5.5 to finish the proof. The formal analysis is stated in Appendix F.2. ∎
Discussion. The transition defined by the condition ${\sf C}\propto{\sf N}^{6}$ identifies a fundamental phase change in the learning trajectory, shifting the regime from one dominated by optimization dynamics to one dominated by statistical scaling. In Stage I (Compute-Starved Stage), the system is compute-starved but capacity-rich. The exponential decay with respect to $\sf C$} suggests that the model is rapidly navigating the loss landscape to find a global minimum, and in this regime, the primary constraint is the duration of training rather than the quality of the data. Conversely, Stage II (Data-Limited Stage) represents the saturation of the optimization process. Once the threshold in Eq. (3) is crossed, the model has effectively ”exhausted” the easily accessible information in the dataset. The shift to a power-law decay of ${\sf C}^{-1/6}$ signifies that further progress requires navigating the high-dimensional geometry of the target function class. This bifurcation explains why empirical scaling laws often only become observable after an initial period of training: Stage I is the transient ”discovery” phase, while Stage II represents the steady-state statistical refinement that defines modern large-scale modeling.
Single-Variable Scaling Law. Moreover, we derive how each variable (training time, dataset size, and model size) affects our result of the scaling law, which we call the single law.
Theorem 7.2 (Single law, informal version of Theorem F.8).
Let all pre-conditions hold as Corollary 6.1. Then, with a probability at least $1-\delta$, there exists:
  • Time-Law. We fix the dataset set $\sf N$} and model size ${\sf M}=\Omega({\sf N}^{3})$, and choosing $\varepsilon=\xi/{\sf N}$, we have $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq\exp(-\varepsilon^{2}\xi^{2}{\sf T}/{\sf N}^{2})+O(\frac{\xi^{2}}{{\sf N}})$.
  • Data-Law. For any dataset size $\sf N$}, we let $\varepsilon=\xi/{\sf N}$, ${\sf M}=\Omega({\sf N}^{3})$ and $T=\widetilde{\Omega}(\frac{{\sf N}^{2}}{\xi})$, we have: $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq O(\frac{\xi^{2}}{{\sf N}})$.
  • Model-Law. We fix the dataset set $\sf N$} and let training time $T=\widetilde{\Omega}(\frac{{\sf M}^{\zeta}}{\xi})$, where $\zeta\in(0,1/3)$. Choosing $\varepsilon=\xi/{\sf M}^{\zeta}$, before the model size $\sf M$} is greater than ${\sf N}^{\frac{3}{2\zeta}}$, we have $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}$ $\Delta{\cal R}(F)\leq\xi^{2}{\sf M}^{-\zeta}$.
Proof Sketch of Theorem 7.2.
This theorem is trivially derived from the intermediate results of Theorem 7.1. For three variables: $\sf T$, $\sf M$, and $\sf N$, we fix two of them to get the single law of the remaining one. The complete derivation is in Appendix F.2. ∎
7.2 Potential Failures of Scaling Law
The Role of the Lambert W Function. The appearance of $W({\sf C}/\xi^{10})$ in Theorem 7.1 is mathematically significant. It indicates that as compute increases, the returns on that compute are not purely power-law distributed but are subject to a logarithmic correction. This implies that at extreme scales, the cost of reducing error becomes slightly more expensive than a simple power law would predict, likely due to the overhead of managing the approximation-generalization gap.
Sensitivity of Noise Level. The term $\xi^{5/3}$ in Theorem 7.1 and the term $\xi^{2}$ of the data-law in Theorem 7.2 indicate that the quality or complexity of the sampled dataset $\mathbb{D}$ drastically shifts the scaling curve. Even with infinite compute, a high-complexity task (large $\xi$) will result in a significantly higher error floor.
The usual experience taught that a larger data size helps neural networks break the information bottleneck to get a better excess risk bound. However, there exists a probability that including more data leads to the noise level $\xi$ growing with the dataset size, we denote it as $\xi({\sf N})$. When $\xi({\sf N})\propto{\sf C}^{1/10}$ or $\xi({\sf N})\propto{\sf N}^{1/2}$, the upper bound on the excess risk degrades to $O(1)$, therefore, such a scaling law would generally become abolished.
The Waste of Model Size Scaling. We review the boundary ${\sf M}<{\sf N}^{\frac{3}{2\zeta}}$ in Theorem 7.2. This suggests that the Model-Law (improving performance by adding parameters) is only effective when the model is small relative to the dataset’s latent complexity. Once the model becomes excessively large (in the highly over-parameterized regime), the risk is limited by the dataset size $\sf N$} rather than the parameter count $\sf M$}.
8 Conclusion
This work presents a comprehensive theoretical framework for rigorously analyzing the scaling law phenomenon in LLMs, specifically addressing the empirically observed power-law relationship between model performance and computational resources from a learning dynamics perspective. By formalizing the dynamics of training sequence-to-sequence multi-layer transformer architectures, we establish a foundational guarantee: under allocation of compute, the generalization excess error of these models converges asymptotically as computational budgets increase with rate i) $\exp({\sf C})$ in the compute-starved stage, and ii) ${\sf C}^{-1/6}$ in the data-limited stage, where $\sf C$} is the computational cost. Furthermore, we quantify the individual influence of three scaling variables (model size, training duration, and dataset scale) on performance. Finally, we identify critical constraints (potential drags) on the reliability of transformer neural scaling, demonstrating that logarithmic corrections from the Lambert W function, the growth of dataset noise levels relative to sample size, and the saturation of model capacity relative to data complexity can lead to diminishing returns or the total breakdown of predictable scaling laws.
Acknowledgement
I would like to sincerely thank Jiangxuan Long, Yiping Lu, Jing Xiong, and Juepeng Zheng for the helpful discussions.
Appendix
Appendix A Technical Preliminary
A.1 Notations
In this paper, we use integer $d$ to denote the dimension of networks. We use $L$ to denote the input length in language models. $\nabla_{x}f(x)$ and $\frac{\mathrm{d}f(x)}{\mathrm{d}x}$ are both means to take the derivative of $f(x)$ with $x$. Let a vector $z\in\mathbb{R}^{n}$. We denote the $\ell_{2}$ norm as $\|z\|_{2}:=(\sum_{i=1}^{n}z_{i}^{2})^{1/2}$, the $\ell_{1}$ norm as $\|z\|_{1}:=\sum_{i=1}^{n}|z_{i}|$, $\|z\|_{0}$ as the number of non-zero entries in $z$, $\|z\|_{\infty}$ as $\max_{i\in[n]}|z_{i}|$. We use $z^{\top}$ to denote the transpose of a $z$. We use $\langle\cdot,\cdot\rangle$ to denote the inner product. Let $A\in\mathbb{R}^{n\times d}$, we use $\operatorname{vec}(A)$ to denote a length $nd$ vector. We denote the Frobenius norm as $\|A\|_{F}:=(\sum_{i\in[n],j\in[d]}A_{i,j}^{2})^{1/2}$. For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. We use $\operatorname*{{\mathbb{E}}}[]$ to denote the expectation. We use $\Pr[]$ to denote the probability. We use $\epsilon$ to denote the error. We define $\lambda_{\min}(\cdot)$ as a function that outputs the minimum eigenvalues of the input matrix, e.g. matrix $A\in\mathbb{R}^{n\times n}$ has eigenvalues $\{\lambda_{1},\lambda_{2},\cdots,\lambda_{n}\}$, $\lambda_{\min}(A)=\min\{\lambda_{1},\lambda_{2},\cdots,\lambda_{n}\}$. For a vector $a\in\mathbb{R}^{n}$, we use $a_{i}$ to denote its $i$-th entry for $i\in[n]$. For a matrix $A\in\mathbb{R}^{n\times d}$, vector $A_{i}\in\mathbb{R}^{d}$ is the $i$-th row for $i\in[n]$ and vector $A_{*,j}\in\mathbb{R}^{n}$ is the $j$-th column for $j\in[d]$. For a function $f:\rightarrow\mathbb{R}^{n}$, we use $f_{i}$ to denote the $i$-th entry of its output. For a function $f:\rightarrow\mathbb{R}^{n\times d}$, we use $f_{i}\in\mathbb{R}^{d}$ to denote the $i$-th row of its output for $i\in[n]$ and us $f_{*,j}\in\mathbb{R}^{n}$ to denote the $j$-th column of its output for $j\in[d]$. We use $\operatorname{\mathbb{I}}\{{\sf E}_{1},{\sf E}_{2},\cdots,{\sf E}_{n}\}$ to denote the indicator for event set $\{{\sf E}_{1},{\sf E}_{2},\cdots,{\sf E}_{n}\}$, only when ${\sf E}_{1},{\sf E}_{2},\cdots,{\sf E}_{n}$ are all true, $\operatorname{\mathbb{I}}\{{\sf E}_{1},{\sf E}_{2},\cdots,{\sf E}_{n}\}=1$; otherwise, it equals to 0. For a vector $a\in\mathbb{R}^{d^{2}}$, function $\operatorname{mat}$ reshapes $a$ to a $d\times d$ matrix, where its $(i,j)$-th entry is $\operatorname{mat}_{i,j}(a)=a_{(i-1)d+j}$ for $(i,j)\in[d]\times[d]$. For a matrix $A\in\mathbb{R}^{n\times d}$, function $\operatorname{vec}$ flattens $A$ to a $nd$-dimensional vector, where its $i$-th entry is $\operatorname{vec}_{i}(A)=A_{\lfloor i/d\rfloor,i-\lfloor i/d\rfloor\cdot d}$ for $i\in[nd]$.
A.2 Original Model Definitions
Definition A.1 (Weights and Initialization).
The weights of the model are denoted as $\theta(t)$ where $t\geq 0$ is time. It contains the weights of each layer $\theta(t)=\{\theta_{(\nu)}(t)\}_{\nu=1}^{N}$, and $\theta_{(\nu)}=\{U_{(\nu)}(t),W_{(\nu)}(t),A_{(\nu)}\}$. For each matrix:
  • Each entry of $U_{(\nu)}(0)\in\mathbb{R}^{d\times d}$ is initialized from the standard Gaussian distribution, formally, $U_{(\nu),k_{1},k_{2}}(0)\sim\mathcal{N}(0,1)$ for any $k_{1},k_{2}\in[d]$.
  • Each entry of $W_{(\nu)}(0)\in\mathbb{R}^{d\times m}$ is initialized from the standard Gaussian distribution, formally, $w_{(\nu),r,k}(0)\sim\mathcal{N}(0,1)$ for any $r\in[m],k\in[d]$, where $w_{(\nu),r}(0)\in\mathbb{R}^{d}$ is the $r$-th column of $W_{(\nu)}(0)$.
  • Each entry of $A_{(\nu)}(0)\in\mathbb{R}^{m\times d}$ is initialized from a $\pm 1$ uniform distribution, formally, $a_{(\nu),r,k}(0)$ $\sim{\sf Uniform}\{-1,+1\}$ for any $r\in[m],k\in[d]$, where $a_{(\nu),r}(0)\in\mathbb{R}^{d}$ is the $r$-th row of $A_{(\nu)}(0)$.
Definition A.2 (Data Distribution).
We denote $F^{*}:{\cal X}\to[C_{1},C_{2}]^{L\times d}$ as the target function. The ${\cal X}\in\mathbb{R}^{L\times d}$ is a combination of $L$ $d$-dimensional balls where $\|X_{\ell}\|_{2}^{2}=\Theta(1),\forall X\in{\cal X},\ell\in[L]$. $C_{1},C_{2}\in\mathbb{R}$ are two fixed constants. The distribution: ${\cal D}=\{(X,F^{*}(X)+\Xi),\text{$\Xi\in\mathbb{R}^{L\times d}$ is some random noise}\}\subset{\cal X}\times{\cal Y}$. The random noise $\Xi\in\mathbb{R}^{L\times d}$ is centered by ${\bf 0}_{L\times d}$ with variance $\max_{\ell\in[L]}\max_{k\in[d]}\operatorname*{\mathrm{Var}}[X_{\ell,k}]=\xi^{2}$.
Definition A.3 (Original Dataset).
The original dataset is $\mathbb{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}\subset{\cal D}$, where $X_{i},Y_{i}=F^{*}(X_{i})+\Xi_{i}\in\mathbb{R}^{L\times d}$. For each data point $(X_{i},Y_{i}),\forall i\in[n]$, it holds that:
  • $\|X_{i,\ell}\|_{2}=\Theta(1)$ for $\ell\in[L]$.
Definition A.4 (Model Functions).
Given an input matrix $X\in\mathbb{R}^{L\times d}$, the model function is given by ($\varepsilon>0$ is the grokking coefficient):
$F(X,\theta(t)):=\varepsilon\cdot F_{(N)}(F_{(N-1)}(\cdots F_{(2)}(F_{(1)}(X+E,\theta(t)),\theta(t))\cdots),\theta(t)).$
We list the original definition of each function as follows:
  • $~F_{(\nu)}(X,\theta(t)):=\frac{\omega}{\sqrt{m}}{\rm ReLU}\left({\rm Softmax}\left(\kappa\cdot XU_{(\nu)}(t)X^{\top}+M\right)XW_{(\nu)}(t)\right)A_{(\nu)}$
  • ${\rm Softmax}\left(A\right):=\operatorname{diag}(\exp(A){\bf 1}_{L})^{-1}\cdot\exp(A)\in\mathbb{R}^{L\times L}$ for $A\in\mathbb{R}^{L\times L}$.
  • ${\rm ReLU}_{\ell,k}(X):=\max\{X_{\ell,k},0\}$ for any $\ell\in[L]$, $k\in[d]$, $X\in\mathbb{R}^{L\times d}$.
  • $M_{\ell_{1},\ell_{2}}:=\begin{cases}0,&~\ell_{1}\geq\ell_{2}\\ -\infty.&~\ell_{1}<\ell_{2}\end{cases},\forall\ell_{1},\ell_{2}\in[L]$.
Definition A.5.
We denote the special notations:
  • We denote $w_{(\nu),r}(t)\in\mathbb{R}^{d}$ as the $r$-th column of $W_{(\nu)}\in\mathbb{R}^{d\times m}$ for $r\in[m]$, $\nu\in[N]$.
  • We denote $a_{(\nu),r}\in\mathbb{R}^{d}$ as the $r$-th row of $A_{(\nu)}\in\mathbb{R}^{m\times d}$ for $r\in[m]$, $\nu\in[N]$.
Definition A.6.
We define the training objective:
${\cal L}(t,\mathbb{D}):=\operatorname*{{\mathbb{E}}}_{(X,Y)\sim\mathbb{D}}[\|F(X,\theta(t))-Y\|_{F}^{2}]$
A.3 Basic Facts and Lemmas
Fact A.7.
For a variable $x\sim\mathcal{N}(0,\sigma^{2})$, then with probability at least $1-\delta$, we have:
$|x|\leq C\sigma\sqrt{\log(1/\delta)}$
Fact A.8.
For an $1$-Lipschitz function $f(\cdot)$, we have:
$|f(x)-f(y)|\leq|x-y|,\forall x,y\in\mathbb{R}^{d}$
Fact A.9.
For a Gaussian variable $x\sim\mathcal{N}(0,\sigma^{2}\cdot I_{d})$ where $\sigma\in\mathbb{R}$, then for any $t>0$, we have:
$\Pr[x\leq t]\leq\frac{2t}{\sqrt{2\pi}\sigma}$
Fact A.10.
For a Gaussian vector $w\sim\mathcal{N}(0,\sigma^{2}\cdot I_{d})$ where $\sigma\in\mathbb{R}$, and a fixed vector $x\in\mathbb{R}^{d}$, we have:
$w^{\top}x\sim\mathcal{N}(0,\sigma^{2}\|x\|_{2}^{2})$
Fact A.11.
For two matrices $H,\widetilde{H}\in\mathbb{R}^{n\times n}$, we have:
$\lambda_{\min}(\widetilde{H})\geq\lambda_{\min}(H)-\|H-\widetilde{H}\|_{F}$
Fact A.12.
The Lipchitz constant of the softmax function is bounded by $O(1)$, such that:
$\|\langle\exp(x),{\bf 1}_{L}\rangle^{-1}\exp(x)-\langle\exp(y),{\bf 1}_{L}\rangle^{-1}\exp(y)\|_{2}\leq O(1)\cdot\|x-y\|_{2},\forall x,y\in\mathbb{R}^{L}.$
Fact A.13.
For a matrix $H\in\mathbb{R}^{n\times n}$, there is $\lambda_{\min}(H\otimes I_{d})=\lambda_{\min}(H)$.
In addition, we state four vital lemmas of concentration inequalities for simplifying analysis:
Lemma A.14 (Hoeffding bound).
Let $X_{1},\cdots,X_{n}$ denote $n$ independent bounded variables in $a_{i},b_{i}]$ for $a_{i},b_{i}\in\mathbb{R}$. Let $X:=\sum_{i=1}^{n}X_{i}$, then we have
$\Pr[|X-\operatorname*{{\mathbb{E}}}[X]|\geq t]\leq 2\exp(-\frac{2t^{2}}{\sum_{i=1}^{n}(b_{i}-a_{i})^{2}})$
Lemma A.15 (Markov’s inequality).
If $X$ is a non-negative random variable and $a>0$, then the probability that $X$ is at least $a$ is at most the expectation of $X$ divided by $a$:
$\Pr[X\geq a]\leq\frac{\operatorname*{{\mathbb{E}}}[X]}{a}$
Lemma A.16 (Chernoff bound).
Let $X=\sum_{i=1}^{n}X_{i}$, where $X_{i}=1$ with probability $p_{i}$ and $X_{i}=0$ with probability $1-p_{i}$, and all $X_{i}$ are independent. Let $\mu=\operatorname*{{\mathbb{E}}}[X]=\sum_{i=1}^{n}p_{i}$. Then
  • $\Pr[X\geq(1+\delta)\mu]\leq\exp(-\delta^{2}\mu/3)$, $\forall\delta>0$;
  • $\Pr[X\leq(1-\delta)\mu]\leq\exp(-\delta^{2}\mu/1)$, $\forall 0<\delta<1$.
Appendix B Gradient Computation and Learning Dynamics
B.1 Simplified Model Definitions
Definition B.1 (Rearranged Dataset).
Given the origin dataset $\mathbb{D}=\{X_{i},Y_{i}\}_{i=1}^{n}\subset\mathbb{R}^{L\times d}\times\mathbb{R}^{L\times d}$. The rearranged dataset is $\mathbb{D}_{\rm rearrange}=\{(X_{i,\leq\ell}+E_{\leq\ell},Y_{i,\ell})\}_{(i,\ell)=(1,1)}^{(n,L)}$, where $X_{i,\leq\ell}+E_{\leq\ell}\in\mathbb{R}^{\ell\times d}$ and $Y_{i,\ell}\in\mathbb{R}^{d}$.
Definition B.2 (Simplified Model Function).
Given the origin dataset $\mathbb{D}=\{X_{i},Y_{i}\}_{i=1}^{n}\subset\mathbb{R}^{L\times d}\times\mathbb{R}^{L\times d}$, we define the compact form of the model function as:
[…]
Here, $i=\lfloor p/L\rfloor$ and $\ell=p{\rm~mod~}L$.
We list the notation-simplified definitions of all functions as follows:
  • (Hidden State) $\Lambda_{(\nu),i}(t):=F_{(\nu)}(\Lambda_{(\nu-1),i}(t),\theta(t))\in\mathbb{R}^{L\times d}$ for $\nu\in[N]$, $\Lambda_{(0),i}(t)=X_{i}+E$.
  • (Attention Scores) $\sigma_{(\nu),(i-1)L+\ell}(X)={\rm Softmax}_{\ell}(\Lambda_{(\nu),i}(t)U_{(\nu)}(t)\Lambda_{(\nu),i}(t)^{\top}+M)\in\mathbb{R}^{L}$.
  • (Attention Output) $o_{(\nu),(i-1)L+\ell}(t):=\Lambda_{(\nu-1),i}(t)^{\top}\cdot\sigma_{(\nu),(i-1)L+\ell}(t)\in\mathbb{R}^{d}$.
  • ($\ell$-th Token of Hidden State) $\mu_{(\nu),(i-1)L+\ell}(t):=\frac{\omega}{\sqrt{m}}\sum_{r=1}^{m}a_{(\nu),r}\cdot\phi(\langle o_{(\nu),(i-1)L+\ell}(t),w_{(\nu),r}(t)\rangle)\in\mathbb{R}^{d}$, where $\phi(x):=\max\{0,x\},\forall x\in\mathbb{R}$. $\mu_{(0),(i-1)L+\ell}(t)=X_{i,\ell}+E_{\ell}$.
  • (Model Output) ${\sf F}_{(i-1)L+\ell}(t)=\varepsilon\cdot\sum_{\nu=0}^{N}\mu_{(\nu),(i-1)L+\ell}(t)\in\mathbb{R}^{d}$.
Lemma B.3.
We have:
${\cal L}(t,\mathbb{D})=\frac{1}{n}\|{\sf F}(t)-{\sf Y}\|_{F}^{2}$
Proof.
Since the decoder-only property of the model function, we have:
$F_{(i-1)L+\ell}(t)=F_{\ell}(X_{i},\theta(t)),\forall(X_{i},Y_{i})\in\mathbb{D},i\in[n],\ell\in[L].$
We then separate each token vector with:
${\cal L}(t,\mathbb{D})=$ $~\operatorname*{{\mathbb{E}}}_{(X,Y)\sim\mathbb{D}}[\|F(X,\theta(t))-Y\|_{F}^{2}]$
$=$ $~\frac{1}{n}\sum_{(X,Y)\in\mathbb{D}}\|F(X,\theta(t))-Y\|_{F}^{2}$
$=$ $~\frac{1}{n}\sum_{(X,Y)\in\mathbb{D}}\sum_{\ell=1}^{L}\|F_{\ell}(X,\theta(t))-Y_{\ell}\|_{2}^{2}$
$=$ $~\frac{1}{n}\|{\sf F}(t)-{\sf Y}\|_{F}^{2},$
where the first three steps follow from simple algebra, and the last step follows from the definitions of $\sf F$} and $\sf Y$}. ∎
B.2 Gradient Computation
Lemma B.4.
For $\nu\in[N]$, we have:
  • Part 1. We have:
    $\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(U_{(\nu)}(t))}=$ $~\frac{\omega\cdot\kappa}{\sqrt{m}}\sum_{p=1}^{nL}(\Lambda_{(\nu-1),i,\ell,*}(t)\otimes\Lambda_{(\nu-1),i}(t))^{\top}\left(\operatorname{diag}(\sigma_{(\nu),p}(t))-\sigma_{(\nu),p}(t)\sigma_{(\nu),p}(t)^{\top}\right)$
    $~\Lambda_{(\nu-1),i}(t)\sum_{r\in[m]}\langle\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu),p}(t)},a_{(\nu),r}\rangle\cdot w_{(\nu),r}(t)\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0\},$
    where $i=\lfloor p/L\rfloor$ and $\ell=p{\rm~mod~}L$.
  • Part 2. For any $r\in[m]$, we have:
    $\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}w_{(\nu),r}(t)}=\frac{\omega}{\sqrt{m}}\sum_{p=1}^{nL}\langle\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu),p}(t)},a_{(\nu),r}\rangle\cdot o_{(\nu),p}(t)\cdot\operatorname{\mathbb{I}}\{w_{(\nu),r}(t)^{\top}o_{(\nu),p}(t)>0\}.$
  • Part 3. For $\nu\in[N]$ and $p\in[nL]$, we have:
    $\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu),p}(t)}=\frac{2\varepsilon}{n}\cdot\left(I_{d}+\operatorname{diag}(G_{(\nu),p}(t))\right)({\sf F}_{p}(t)-{\sf Y}_{p}),$
    where
    $G_{(\nu),p}(t)=$ $~\frac{\omega}{\sqrt{m}}\Big(\sigma_{(\nu),p,\ell}(t)\cdot I_{d}$
    $~+\kappa U_{(\nu)}(t)\Lambda_{(\nu-1),i}(t)^{\top}(\operatorname{diag}({\bf 1}_{n}-\frac{1}{2}e_{\ell}))\cdot\left(\operatorname{diag}(\sigma_{(\nu),p}(t))-\sigma_{(\nu),p}(t)\sigma_{(\nu),p}(t)^{\top}\right)\Lambda_{(\nu-1),i}(t)\Big)$
    $~\cdot\sum_{r\in[m]}\langle\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu+1),p}(t)},a_{(\nu),r}\rangle\cdot w_{(\nu),r}(t)\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0\}),$
    and $G_{(N),p}(t)={\bf 0}_{d}$.
    Here, we denote $i=\lfloor p/L\rfloor$ and $\ell=p{\rm~mod~}L$. $e_{\ell}$ is a $n$-dimensional one-hot vector with the $r$-th entry equal $1$.
Proof.
We omit the lemma proof since the gradient computation trivially follows from basic algebra and multivariable chain rules. ∎
B.3 Learning Dynamics
We give the definition of NTK:
Definition B.5.
We define the kernel matrix at $\nu$-th layer as $H_{(\nu)}\in\mathbb{R}^{nL\times nL}$ and its $(i,j)$-th entry ($\forall(p,q)\in[nL]\times[nL]$) is defined as:
$H_{(\nu),p,q}(t):=\underbrace{\langle\beta_{(\nu),p}(t),\beta_{(\nu),q}(t)\rangle}_{\text{kernel w.r.t. $W_{(\nu)}(t)$}}+\underbrace{\langle\gamma_{(\nu),p}(t),\gamma_{(\nu),q}(t)\rangle}_{\text{kernel w.r.t. $U_{(\nu)}(t)$}},$
Here, we let:
$\beta_{(\nu),p}(t):=$ $~\frac{\omega}{\sqrt{m}}\underbrace{o_{(\nu),p}(t)}_{d\times 1}\otimes\underbrace{{\bf 1}_{W_{(\nu)}(t)^{\top}o_{(\nu),p}(t)>0}}_{m\times 1}\in\mathbb{R}^{md},$
$\gamma_{(\nu),p}(t):=$ $~\frac{\omega\cdot\kappa}{\sqrt{m}}\underbrace{(\Lambda_{(\nu-1),i,\ell,*}(t)\otimes\Lambda_{(\nu-1),i}(t))^{\top}}_{d^{2}\times L}\underbrace{\left(\operatorname{diag}(\sigma_{(\nu),p}(t))-\sigma_{(\nu),p}(t)\sigma_{(\nu),p}(t)^{\top}\right)}_{L\times L}$
$~\underbrace{\Lambda_{(\nu-1),i}(t)}_{L\times d}\sum_{r\in[m]}\underbrace{w_{(\nu),r}(t)\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0\}}_{d\times 1}\in\mathbb{R}^{d^{2}},$
where $\otimes$ is the Kronecker product and $i=\lfloor p/L\rfloor$, $\ell=p{~\rm mod~}L$. The indicator vector ${\bf 1}_{W_{(\nu)}(t)^{\top}o_{(\nu),p}(t)>0}\in\{0,1\}^{m}$ where its $r$-th entry is $\operatorname{\mathbb{I}}\{\left(W_{(\nu)}(t)^{\top}o_{(\nu),p}(t)\right)_{r}>0\}$ for $r\in[m]$.
We restate Lemma 4.1 below as its formal version:
Lemma B.6 (Formal version of Lemma 4.1).
The learning dynamics of the multi-layer transformer Eq. (1) is given by:
$\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]=-\sum_{\nu\in[N]}{\underbrace{\operatorname{vec}\left(\frac{\mathrm{d}}{\mathrm{d}\mu_{(\nu)}(t)}{\cal L}(t,\mathbb{D})\right)^{\top}}_{1\times nLd}}\cdot\underbrace{\left(H_{(\nu)}(t)\otimes I_{d}\right)}_{nLd\times nLd}\cdot\underbrace{\operatorname{vec}\left(\frac{\mathrm{d}}{\mathrm{d}\mu_{(\nu)}(t)}{\cal L}(t,\mathbb{D})\right)}_{nLd\times 1}$
where $\mu_{(\nu)}(t)$ is a $nL\times d$ matrix, $\mu_{(\nu),p}(t)$ is the $(p\text{ mod }L)$-th row of $v$-th layer output regarding to input matrix $X_{\lfloor p/L\rfloor}$ for any $p\in[nL]$ and $\nu\in[N]$.
Proof.
We have:
$\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]=$ $~\operatorname*{{\mathbb{E}}}[\sum_{\nu=1}^{N}(\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(W_{(\nu)}(t))}^{\top}\frac{\mathrm{d}\operatorname{vec}(W_{(\nu)}(t))}{\mathrm{d}t}+\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(U_{(\nu)}(t))}^{\top}\frac{\mathrm{d}\operatorname{vec}(U_{(\nu)}(t))}{\mathrm{d}t})]$
$=$ $~-\sum_{\nu=1}^{N}(\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(W_{(\nu)}(t))}^{\top}\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(W_{(\nu)}(t))}+\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(U_{(\nu)}(t))}^{\top}\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\operatorname{vec}(U_{(\nu)}(t))})$
$=$ $~-\sum_{\nu=1}^{N}\sum_{p=1}^{nL}\sum_{q=1}^{nL}\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu),p}(t)}^{\top}\left(\langle\beta_{(\nu),p}(t),\beta_{(\nu),q}(t)\rangle+\langle\gamma_{(\nu),p}(t),\gamma_{(\nu),q}(t)\rangle\right)\cdot I_{d}\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu),q}(t)}$
$=$ $~-\sum_{\nu\in[N]}{\operatorname{vec}\left(\frac{\mathrm{d}}{\mathrm{d}\mu_{(\nu)}(t)}{\cal L}(t,\mathbb{D})\right)^{\top}}\cdot\left(H_{(\nu)}(t)\otimes I_{d}\right)\cdot\operatorname{vec}\left(\frac{\mathrm{d}}{\mathrm{d}\mu_{(\nu)}(t)}{\cal L}(t,\mathbb{D})\right),$
where the first step follows from the chain rule, the second step follows from the gradient flow (Eq. (2)), the third step follows from Part 1, Part 2 and Part 3 of Lemma B.4 and the definitions of $\beta_{(\nu),p}(t)$ and $\gamma_{(\nu),p}(t)$, the last step follows from the definition of $H_{(\nu)}(t)$. ∎
Appendix C Toolkit: Helpful Boundaries
We give the lemma about all help bounds in the range of this paper as follows:
Lemma C.1.
Denote failure probability $\delta\in(0,0.1)$. Define $B:=\max\{O(\sqrt{\log(Lmd/\delta)}),1\}$. Assuming there exists a constant $R\in(0,1)$ satisfying $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$ for $r\in[m]$ and $\nu\in[N]$ and $R\in(0,1)$.
We mark the indices as: $r\in[m]$, $p\in[nL]$, $\ell_{1},\ell_{2}\in[L]$, and $\nu\in[N]$. We denote $i=\lfloor p/L\rfloor$, $\ell=p{\rm~mod~}L$, $\ell^{\prime}\in[\ell]$.
If Definition 5.2 holds, then with a probability at least $1-\delta$, we have:
  • Basic Bounds.
    • Part 1. $\|w_{(\nu),r}(0)\|_{2}\leq O(\sqrt{d}B)$.
    • Part 2. $\|U_{(\nu)}(0)\|_{F}\leq O(dB)$.
    • Part 3. $\|w_{(\nu),r}(t)\|_{2}\leq O(\sqrt{d}B)$.
    • Part 4. $\|U_{(\nu)}(t)\|_{F}\leq O(dB)$.
    • Part 5. $\|\Lambda_{(\nu),i,\ell_{1}}(t)\|_{2}=\Theta(1)$.
    • Part 6. $\|\Lambda_{(\nu),i}(t)U_{(\nu)}(t)\Lambda_{(\nu),i}(t)^{\top}\|_{\infty}\leq O(dB)$.
    • Part 7. $\exp_{\ell_{1},\ell_{2}}(\Lambda_{(\nu),i}(t)U_{(\nu)}(t)\Lambda_{(\nu),i}(t)^{\top})\in[\exp(-O(dB),\exp(O(dB))]$.
    • Part 8. For $\ell^{\prime}\in[\ell]$, $\sigma_{(\nu),p,\ell^{\prime}}(t)\in[\exp(-O(dB))/L,1]$.
  • Perturbation Bounds.
    • Part 9. $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$.
    • Part 10. $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$.
    • Part 11. $\|\Lambda_{(\nu),i,\ell}(t)-\Lambda_{(\nu),i,\ell}(0)\|_{2}\leq O(R)$.
    • Part 12. $\|\sigma_{(\nu),p}(t)-\sigma_{(\nu),p}(0)\|_{2}\leq O(\sqrt{L}R)$.
    • Part 13. $\|o_{(\nu),p}(t)-o_{(\nu),p}(0)\|_{2}\leq O(\sqrt{L}R)$.
  • Gradient and Function Norms.
    • Part 14. ${\cal L}(t,\mathbb{D})\leq O(Ld)$.
    • Part 15. $\|\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu)}(t)}\|_{F}^{2}\asymp\varepsilon^{2}\cdot{\cal L}(t,\mathbb{D})$.
    • Part 16. $\|\gamma_{(\nu),p}(t)\|_{2}\leq o(1/\sqrt{m})$.
Proof.
Proof of Part 1. This proof follows from the initialization of $W_{(\nu)}(0)$ and the union bound of the tail bound of the Gaussian distribution (Fact A.7).
Proof of Part 2. This proof follows from initialization of $U_{(\nu)}(0)$ and the union bound of the tail bound of the Gaussian distribution (Fact A.7).
Proof of Part 3. This proof combines $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$, triangle inequality and $R\leq B$.
Proof of Part 4. This proof combines $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$, triangle inequality and $R\leq B$.
Proof of Part 5. We have:
$\|\Lambda_{(0),i,\ell}(t)\|_{2}=\Theta(1).$
Therefore, we have:
$\|o_{(1),p}(t)\|_{2}=\|\Lambda_{(0),i}(t)^{\top}\sigma_{(\nu),p}(t)\|_{2}=\Theta(1),$
where this step follows from the property that the sum of any softmax vector is $1$.
Therefore, we can show that:
$\|\mu_{(\nu),p}(t)\|_{2}=$ $~\|\frac{\omega}{\sqrt{m}}\sum_{r=1}^{m}a_{(1),r}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)\|_{2}$
$\leq$ $~\frac{\omega\sqrt{d}}{\sqrt{m}}\|\sum_{r=1}^{m}a_{(1),r}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)\|_{\infty}$
$\leq$ $~\frac{\omega\sqrt{d}}{\sqrt{m}}\max_{k\in[d]}|\sum_{r=1}^{m}a_{(1),r,k}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)|$
where the first step follows from the definition of $\mu_{(\nu),p}(t)$, the second step follows from simple algebras, the third step follows from the definition of infinite norm.
We apply Hoeffding bound to each variable $a_{(1),r,k}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)$, we have:
$~|a_{(1),r,k}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)|\leq O(\sqrt{d}B),$
$~\operatorname*{{\mathbb{E}}}[a_{(1),r,k}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)]=0.$
With a probability at least $1-\delta$, we have:
$\|\mu_{(\nu),p}(t)\|_{2}\leq$ $~\frac{\omega\sqrt{d}}{\sqrt{m}}\max_{k\in[d]}|\sum_{r=1}^{m}a_{(1),r,k}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)|$
$\leq$ $~\frac{\omega\sqrt{d}}{\sqrt{m}}\cdot O(\sqrt{md}B)\sqrt{\log(m/\delta)}\leq O(\omega dB^{2}).$
Hence,
$\|\Lambda_{(1),i,\ell}(t)\|_{2}=$ $~\|\frac{\omega}{\sqrt{m}}\sum_{r=1}^{m}a_{(1),r}\cdot\phi(\langle o_{(1),p}(t),w_{(1),r}(t)\rangle)+\Lambda_{(0),i,\ell}(t)\|_{2}$
$=$ $~\Theta(1)\pm O(\omega dB^{2})$
$=$ $~\Theta(1)\pm o(\frac{1}{N})$
the last step follows from choosing $\omega=o(\frac{1}{NdB^{2}})$.
By induction, we can get:
$\|\Lambda_{(\nu),i,\ell}(t)\|_{2}=\Theta(1),$
and
$\|o_{(\nu),p}(t)\|_{2}=\Theta(1).$ (4)
Proof of Part 6. We have:
$\|\Lambda_{(\nu),i}(t)U_{(\nu)}(t)\Lambda_{(\nu),i}(t)^{\top}\|_{\infty}=$ $~\max_{(\ell_{1},\ell_{2})\in[L]\times[L]}\left|\Lambda_{(\nu),i,\ell_{1}}(t)^{\top}U_{(\nu)}(t)\Lambda_{(\nu),i,\ell_{2}}(t)\right|$
$\leq$ $~O(dB)$
where the first step follows from simple algebra, the second step follows from Part 4 and Part 5 of this lemma and Cauchy-Schwarz inequality.
Proof of Part 7. This proof follows from Part 3 of this lemma and simple algebra.
Proof of Part 8. Following Part 7 of this lemma, we can show that:
$\sum_{\ell^{\prime}\in[\ell]}\exp(\Lambda_{(\nu),i,\ell}(t)^{\top}U_{(\nu)}(t)\Lambda_{(\nu),i,\ell^{\prime}}(t))\leq\ell\exp(O(dB))\leq L\exp(O(dB)).$
Then we can show that:
$\sigma_{(\nu),p,\ell^{\prime}}(t)\geq$ $~\exp(\Lambda_{(\nu),i,\ell}(t)^{\top}U_{(\nu)}(t)\Lambda_{(\nu),i,\ell^{\prime}}(t))/\sum_{\ell^{\prime}\in[\ell]}\exp(\Lambda_{(\nu),i,\ell}(t)^{\top}U_{(\nu)}(t)\Lambda_{(\nu),i,\ell^{\prime}}(t))$
$\geq$ $~\frac{\exp(-O(dB))}{L\exp(O(dB))}$
$\geq$ $~\frac{\exp(-O(dB))}{L}.$
Combining the previous results, we obtain the result of this part.
Proof of Part 9. Directly from the lemma condition $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{F}\leq R$.
Proof of Part 10. Directly from the lemma condition $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$.
Proof of Part 11, 12 and 13. When $\nu=1$, we have:
$~\|\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i,\ell}(0)\|_{F}$
$=$ $~\sqrt{\sum_{k\in[d]}\left(\frac{\omega}{\sqrt{m}}\cdot\sum_{r=1}^{m}a_{(1),r,k}\left(\phi(\langle w_{(1),r}(t),o_{(1),p}(t)\rangle)-\phi(\langle w_{(1),r}(0),o_{(1),p}(0)\rangle)\right)\right)^{2}},$
besides, $\Lambda_{(0),i,\ell}(t)=\Lambda_{(0),i,\ell}(0)$.
Next, we have:
$\|\sigma_{(1),p}(t)-\sigma_{(1),p}(0)\|_{2}\leq$ $~O(1)\cdot\|\Lambda_{(0),i}(0)(U_{(1)}(t)-U_{(1)}(0))^{\top}\Lambda_{(0),i,\ell}(0)\|_{2}$
$\leq$ $~O(\sqrt{L}R)$
where the first step follows from the definition of $\sigma_{(1),i}(t)$ and Fact A.12, the second step follows from $\|U_{(1)}(t)-U_{(1)}(0)\|_{F}\leq R$, $\|\Lambda_{(0),i}(0)\|_{F}=O(\sqrt{L})$, $\|\Lambda_{(0),i,\ell}(0)\|_{2}=O(1)$ and Cauchy-Schwarz inequality.
Thus, we can show that
$\|o_{(1),i}(t)-o_{(1),i}(0)\|_{2}=$ $~\|\Lambda_{(0),i}(0)^{\top}\sigma_{(1),p}(t)-\Lambda_{(0),i}(0)^{\top}\sigma_{(1),p}(0)\|_{2}$
$=$ $~\sqrt{\sum_{\ell^{\prime}=1}^{\ell}(\sigma_{(1),p,\ell^{\prime}}(t)-\sigma_{(1),p,\ell^{\prime}}(t))^{2}\|\Lambda_{(0),i,\ell}(0)\|_{2}^{2}}$
$\leq$ $~\sqrt{\sum_{\ell^{\prime}=1}^{\ell}(\sigma_{(1),p,\ell^{\prime}}(t)-\sigma_{(1),p,\ell^{\prime}}(t))^{2}O(1)}$
$=$ $~O(1)\cdot\|\sigma_{(1),p}(t)-\sigma_{(1),p}(t)\|_{2}$
$\leq$ $~O(\sqrt{L}R),$
where these steps follow from some basic algebra, Fact A.12 and Part 10 of this Lemma.
For a certain index $k\in[d]$, we apply Hoeffding’s inequality (Lemma A.14) to each random variable $a_{(1),r}\cdot\phi(\langle w_{(1),t}(t),o_{(1),i}(t)\rangle)$, besides, we have:
$~|a_{(1),r,k}\cdot(\phi(\langle w_{(1),r}(t),o_{(1),i}(t)\rangle)-\phi(\langle w_{(1),r}(0),o_{(1),i}(0)\rangle))|$
$\leq$ $~\max\{\|\|w_{(1),r}(0)\|\cdot O(\sqrt{L}R),\|o_{(1),i}(t)\|_{2}\cdot O(R)\}$
$=$ $~O(\sqrt{Ld}BR),$
$~\operatorname*{{\mathbb{E}}}[a_{(1),r,k}\cdot(\phi(\langle w_{(1),r}(t),o_{(1),i}(t)\rangle)-\phi(\langle w_{(1),r}(0),o_{(1),i}(0)\rangle))]=0.$
Then with a probability at least $1-\delta$, we have:
$\frac{\omega}{\sqrt{m}}\sum_{r=1}^{m}a_{(1),r,k}\cdot\phi(\langle w_{(1),t}(t),o_{(1),i}(t)\rangle)\leq O(\frac{\omega}{\sqrt{m}}BR)\sqrt{Lmd\log(1/\delta)}\leq O(R/(Bd\sqrt{Ld}))$
where the last step follows from choosing $\omega=o(\frac{1}{\sqrt{L}d^{2}B^{3}})$.
We therefore get $\|\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i,\ell}(0)\|_{2}\leq O(R)$ for all $\ell\in[\ell]$ and $\|\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i,\ell}(0)\|_{\infty}\leq O(R/\sqrt{d})$.
When $\nu=2$, we have:
$~\|\sigma_{(2),p}(t)-\sigma_{(2),p}(t)\|_{2}$
$\leq$ $~O(1)\cdot\|\Lambda_{(1),i}(t)U_{(2)}(t)^{\top}\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i}(0)U_{(2)}(0)^{\top}\Lambda_{(1),i,\ell}(0)\|_{2}$
$=$ $~O(1)\cdot\|(\Lambda_{(1),i}(t)\otimes\Lambda_{(1),i,\ell}(t))\operatorname{vec}(U_{(2)}(t))\rangle-(\Lambda_{(1),i}(t)\otimes\Lambda_{(1),i,\ell}(t))\operatorname{vec}(U_{(2)}(0))\|_{2}$
$\leq$ $~O(1)\cdot\max\{\|\Lambda_{(1),i}(t)\otimes\Lambda_{(1),i,\ell}(t)\|_{F}\cdot O(R),$
$~\quad\quad~~~~~~~~~~~~\|U_{(2)}(0)\|_{F}\cdot\|\Lambda_{(1),i}(t)\otimes\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i}(0)\otimes\Lambda_{(1),i,\ell}(0)\|_{F}\}$
$\leq$ $~O(\sqrt{L}R)$
where the first step follows from simple algebra and Fact A.12, the second step follows from a basic tensor trick, the third step follows from simple algebra, triangle inequality, Part 4 of this lemma and $\|\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i,\ell}(0)\|_{\infty}\leq O(R/(Bd\sqrt{Ld}))$.
Hence, we have:
$\|o_{(2),i}(t)-o_{(2),i}(0)\|_{2}=$ $~\|\Lambda_{(1),i}(t)^{\top}\sigma_{(2),p}(t)-\Lambda_{(1),i}(0)^{\top}\sigma_{(2),p}(0)\|_{2}$
$\leq$ $~\max\{\|\Lambda_{(1),i}(t)\|\cdot O(RLdB),O(1)\cdot O(\sqrt{L}R)\}$
$\leq$ $~O(\sqrt{L}R),$
where these steps follow from some basic algebra, Fact A.12 and Part 10 of this Lemma.
Here, we similarly apply Hoeffding inequality (Lemma A.14), and get: with a probability at least $1-\delta$, we have:
$\|\Lambda_{(2),i,\ell}(t)-\Lambda_{(2),i,\ell}(0)\|_{2}\leq O(R).$
By induction, we obtain the results of Part 11, 12, and 13.
Proof of Part 14. We have:
${\cal L}(t,\mathbb{D})=$ $~\frac{1}{n}\|{\sf F}(t)-{\sf Y}\|_{F}^{2}$
$\leq$ $~L\max_{p\in[nL]}\|{\sf F}_{p}(t)-{\sf Y}_{p}\|_{2}^{2}$
$\leq$ $~L\max_{p\in[nL]}\left(\|{\sf F}_{p}(t)\|_{2}+\|{\sf Y}_{p}\|_{2}\right)^{2}$
$\leq$ $~L\left(O(1)+O(\sqrt{d})\right)^{2}$
$\leq$ $~O(Ld),$
where the first step follows from Lemma B.3, the second step follows from simple algebras, the third step follows from the Cauchy-Schwartz inequality, the last two steps follow from Definition A.3 and simple algebras.
Proof of Part 15. It is easy to prove:
$\|\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(N)}(t)}\|_{F}\leq O(\sqrt{d}).$
Besides, we can obtain:
$~\kappa U_{(N)}(t)\Lambda_{(N-1),i}(t)^{\top}(\operatorname{diag}({\bf 1}_{n}-\frac{1}{2}e_{\ell}))\cdot\left(\operatorname{diag}(\sigma_{(N),p}(t))-\sigma_{(N),p}(t)\sigma_{(N),p}(t)^{\top}\right)\Lambda_{(N-1),i}(t)$
$\leq$ $~O(\kappa L^{2}dB)$
$\leq$ $~O(L^{2}dB).$
Then, by Hoeffding inequality (Lemma A.14), we have:
$\|G_{(N-1),p}(t)\|_{2}^{2}\leq\omega\cdot O(L^{2}d^{2.5}B^{3}).$
Choosing $\omega=o(\frac{1}{L^{2}d^{2.5}B^{3}})$, we have: $\|G_{(N-1),p}(t)\|_{2}^{2}\leq o(1)$.
By induction, we can show that: $\|\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu)}(t)}\|_{F}\leq O(\sqrt{d}).$
Similarly, we have:
$\|\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}\mu_{(\nu)}(t)}\|_{F}^{2}\asymp\varepsilon^{2}\cdot{\cal L}(t,\mathbb{D}).$
Proof of Part 16. This proof is trivially similar with bounding $\|G_{(\nu),p}(t)\|_{2}^{2}\leq o(1)$, then we choose $\kappa=1/\sqrt{m}$ to meet the result.
This completes the proof of all parts of the lemma. ∎
Appendix D Kernel Perturbation
Here is a formal version of confirmation of Lemma 5.3:
Lemma D.1 (Formal version of Lemma 5.3).
Assuming Assumption 5.1 and Definition 5.2 hold, denote the failure probability $\delta\in(0,0.1)$, then the kernel perturbation bound is:
$\Pr\left[\lambda_{\min}(H_{(\nu)}(t))<\lambda/2\right]<\delta.$
Therefore, bounding loss dynamics is given by ($C>0$ is some constant):
$\Pr\left[\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]>-C\cdot\omega\lambda N\cdot{\cal L}(t,\mathbb{D})\right]<\delta.$
Proof.
Proof of Part 1. First, we assume the condition $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$ as Lemma C.1, which will be confirmed as a provable property in the further analysis.
For any $p,q\in[nL]$, we have:
$~|H_{(\nu),p,q}^{\prime}(t)-H_{(\nu),p,q}^{\prime}(0)|$
$=$ $~\frac{1}{m}|o_{(\nu),p}(t)^{\top}o_{(\nu),p}(t)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0,o_{(\nu),q}(t)^{\top}w_{(\nu),r}(t)>0\}$
$~-o_{(\nu),p}(0)^{\top}o_{(\nu),p}(0)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(0)^{\top}w_{(\nu),r}(0)>0,o_{(\nu),q}(0)^{\top}w_{(\nu),r}(0)>0\}|$
$\leq$ $~\frac{1}{m}(Q_{(\nu),p,q,1}+Q_{(\nu),p,q,2}+Q_{(\nu),p,q,3}),$
where the first step follows from the definition of $H_{(\nu),p,q}^{\prime}(t)$, and the second step follows from the triangle inequality and defining:
$Q_{(\nu),p,q,1}:=$ $~|o_{(\nu),p}(t)^{\top}o_{(\nu),p}(t)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0,o_{(\nu),q}(t)^{\top}w_{(\nu),r}(t)>0\}$
$~-o_{(\nu),p}(0)^{\top}o_{(\nu),p}(t)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0,o_{(\nu),q}(t)^{\top}w_{(\nu),r}(t)>0\}|,$
$Q_{(\nu),p,q,2}:=$ $~|o_{(\nu),p}(0)^{\top}o_{(\nu),p}(t)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0,o_{(\nu),q}(t)^{\top}w_{(\nu),r}(t)>0\}$
$~-o_{(\nu),p}(0)^{\top}o_{(\nu),p}(0)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0,o_{(\nu),q}(t)^{\top}w_{(\nu),r}(t)>0\}|,$
$Q_{(\nu),p,q,3}:=$ $~|o_{(\nu),p}(0)^{\top}o_{(\nu),p}(0)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(t)^{\top}w_{(\nu),r}(t)>0,o_{(\nu),q}(t)^{\top}w_{(\nu),r}(t)>0\}$
$~-o_{(\nu),p}(0)^{\top}o_{(\nu),p}(0)\sum_{r\in[m]}\operatorname{\mathbb{I}}\{o_{(\nu),p}(0)^{\top}w_{(\nu),r}(0)>0,o_{(\nu),q}(0)^{\top}w_{(\nu),r}(0)>0\}|.$
We assume a constant $R$ that satisfies $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$.
Bounding $Q_{(\nu),p,q,1}$. We have:
$Q_{(\nu),p,q,1}\leq$ $~\Big|\Big(o_{(\nu),p}(t)-o_{(\nu),p}(0)\Big)^{\top}o_{(\nu),q}(t)$
$~\cdot\sum_{r=1}^{m}\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}\Big|$
$\leq$ $~m\Big|\Big(o_{(\nu),p}(t)-o_{(\nu),p}(0)\Big)^{\top}o_{(\nu),q}(t)\Big|$
$\leq$ $~m\|o_{(\nu),p}(t)-o_{(\nu),p}(0)\|_{2}\cdot\|o_{(\nu),q}(t)\|$
$\leq$ $~m\cdot O(\sqrt{L}R),$
where the first step follows from the definition of $Q_{(\nu),p,q,1}$ and Cauchy-Schwarz inequality, the second step follows from $\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}\leq 1$, the third step follows from Cauchy-Schwarz inequality, the last step follows from Part 6 and Part 13 of Lemma C.1 and $\|o_{(\nu),q}(t)\|\leq O(1)$.
Bounding $Q_{(\nu),p,q,2}$. We omit this proof since it is similar to the proof of bounding $Q_{(\nu),p,q,1}$, we have:
$Q_{(\nu),p,q,2}\leq m\cdot O(\sqrt{L}R).$
Bounding $Q_{(\nu),p,q,3}$. We define the following event:
${\sf E}_{(\nu),p,r}:=$ $~\{\exists w\in\mathbb{R}^{d}:\|w-w_{(\nu),r}(0)\|_{2}\leq R,$
$~\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle>0\}\neq\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0\}\}.$
It is easy to hold that, once:
$~|\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle|\geq O(\sqrt{L}R)\cdot O(\sqrt{d}B)+O(1)\cdot R$
$\iff$ $~|\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle|\geq O(\sqrt{Ld}BR),$
the event ${\sf E}_{(\nu),p,r}$ is false, since we combining Part 1, Part 9 and Part 13 of Lemma C.1, $\|o_{(\nu),q}(t)\|\leq O(1)$ and some simple algebra.
Following Fact A.10, we have:
$\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle\sim\mathcal{N}(0,\|o_{(\nu),p}(0)\|_{2}^{2}).$
We have:
$\|o_{(\nu),p}(0)\|_{2}^{2}=\sum_{\ell^{\prime}=1}^{\ell}\sigma_{(\nu),p,\ell^{\prime}}(t)^{2}\|\Lambda_{(\nu),p,\ell^{\prime}}(t)\|_{2}^{2}=O(1)\cdot\sum_{\ell=1}^{L_{i}}\sigma_{(\nu),p,\ell^{\prime}}(t)^{2}\geq\ell\exp(-O(dB))\geq\exp(-O(dB)),$
where $\ell=p{\rm~mod~}L$.
Then, the anti-concentration of $\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle$ shows that:
$\Pr[{\sf E}_{(\nu),p,r}]\leq$ $~\frac{1}{\Theta(1)\cdot\|o_{(\nu),p}(0)\|_{2}}O(\sqrt{Ld}BR)$
$\leq$ $~O(\sqrt{L}R)\cdot\exp(O(dB)),$
where the first step follows from Fact A.9, the second step follows from simple algebra and $\|o_{(\nu),p}(0)\|_{2}^{2}\geq\exp(-O(dB))$.
We have:
$\operatorname*{{\mathbb{E}}}[Q_{(\nu),p,q,3}]=$ $~\operatorname*{{\mathbb{E}}}\Big[\Big|o_{(\nu),p}(0)^{\top}o_{(\nu),q}(0)\Big|$
$~\cdot\sum_{r=1}^{m}\Big|\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}$
$~\quad\quad~~-\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle>0,\langle o_{(\nu),q}(0),w_{(\nu),r}(0)\rangle>0\}\Big|\Big]$
$\leq$ $~\operatorname*{{\mathbb{E}}}\Big[O(1)\cdot\sum_{r=1}^{m}\Big|\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}$
$~-\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle>0,\langle o_{(\nu),q}(0),w_{(\nu),r}(0)\rangle>0\}\Big|\Big]$
$\leq$ $~O(1)\cdot\sum_{r=1}^{m}\operatorname*{{\mathbb{E}}}\Big[\Big|\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}$
$~-\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(0),w_{(\nu),r}(0)\rangle>0,\langle o_{(\nu),q}(0),w_{(\nu),r}(0)\rangle>0\}\Big|\Big]$
$\leq$ $~O(1)\cdot\sum_{r=1}^{m}\operatorname*{{\mathbb{E}}}\Big[\operatorname{\mathbb{I}}\{{\sf E}_{(\nu),p,r}\cup{\sf E}_{(\nu),q,r}\}\Big]$
$\leq$ $~O(1)\cdot\sum_{r=1}^{m}\exp(O(dB))\cdot RL^{\nu/2}$
$\leq$ $~mO(\sqrt{L}R)\cdot\exp(O(dB)),$
where the first step follows from the definition of $Q_{(\nu),p,q,3}$, the second step follows from $\|o_{(\nu),p}(0)\|_{2}\leq O(1)$, the third and fourth step follow from the rules of expectation, the last two steps follow from $\Pr[{\sf E}_{(\nu),p,r}]\leq O(\sqrt{L}R)\cdot\exp(O(dB))$ and simple algebra.
Hence, using Markov’s inequality (Lemma A.15), with a probability at least $1-\delta$, we have:
$Q_{(\nu),p,q,3}\leq m\cdot O(\sqrt{L}R)\cdot\exp(O(dB))/\delta.$
Combine the upper bounds on three terms, we have:
$|H_{(\nu),p,q}^{\prime}(t)-H_{(\nu),p,q}^{\prime}(0)|\leq O(\sqrt{L}R)\cdot\exp(O(dB))/\delta.$
We have:
$\|H_{(\nu)}^{\prime}(t)-H_{(\nu)}^{\prime}(0)\|_{F}=$ $~\sqrt{\sum_{p=1}^{nL}\sum_{q=1}^{nL}(H_{(\nu),p,q}(t)-H_{(\nu),p,q}(0))^{2}}$
$\leq$ $~O(nL^{1.5}R)\cdot\exp(O(dB))/\delta.$ (5)
Following Fact A.11 and choose
$R\leq\frac{\omega\lambda\delta}{nL^{1.5}\exp(O(dB))},$ (6)
we have:
$\lambda_{\min}(H_{(\nu)}^{\prime}(t))\geq\frac{3}{4}\omega\lambda.$
It is trivial to have (Part 16 of Lemma C.1):
$\|\gamma_{(\nu),p}(t)\|_{2}\leq o(\frac{1}{\sqrt{m}}).$
Therefore, we have:
$|H_{(\nu),p,q}(t)-H_{(\nu),p,q}^{\prime}(t)|=$ $~|\langle\gamma_{(\nu),p}(t),\gamma_{(\nu),q}(t)\rangle|$
$\leq$ $~o(1/m).$
Furthermore, we have:
$\|H_{(\nu)}(t)-H_{(\nu)}^{\prime}(t)\|_{F}\leq o(\frac{n}{m})\leq\omega\lambda/4,$
where the last step follows from $m=\Omega(n/(\omega\lambda))$.
Similarly, following Fact A.11 and combining the previous result, we have:
$\lambda_{\min}(H_{(\nu)}(t))\geq\omega\lambda/2.$
Following Fact A.13, we have:
$\lambda_{\min}(H_{(\nu)}(t)\otimes I_{d})\geq\omega\lambda/2.$
Combining Lemma B.6, Part 15 of Lemma C.1 and Fact A.11, we can show that:
$\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]\leq-C\cdot\omega\lambda N\cdot\varepsilon^{2}\cdot{\cal L}(t,\mathbb{D}).$
Solving this ODE, we have:
$\operatorname*{{\mathbb{E}}}[{\cal L}(T,\mathbb{D})]\leq\exp(-C\cdot\varepsilon^{2}\omega\lambda NT)\cdot{\cal L}(0,\mathbb{D}).$ (7)
Bounding Gradient Norm. Following Part 3 of Lemma C.1, we have:
$~\max_{T\geq 0}\max_{\nu\in[N]}\max_{r\in[m]}\|w_{(\nu),r}(T)-w_{(\nu),r}(0)\|_{2}$
$\leq$ $~\max_{T\geq 0}\int_{0}^{T}\max_{\nu\in[N]}\max_{r\in[m]}\|\frac{\mathrm{d}{\cal L}(s,\mathbb{D})}{\mathrm{d}w_{(\nu),r}(s)}\|_{2}\mathrm{d}s$
$\leq$ $~\max_{T\geq 0}\int_{0}^{T}\exp(-C\cdot\varepsilon^{2}\omega\lambda Ns)\cdot\varepsilon^{2}{\cal L}(0,\mathbb{D})\cdot\frac{\omega}{\sqrt{m}}\cdot O(\sqrt{d}B)\mathrm{d}s$
$\leq$ $~\max_{T\geq 0}\int_{0}^{T}\exp(-C\cdot\varepsilon^{2}\omega\lambda Ns)\cdot\frac{\varepsilon^{2}}{\sqrt{m}}\mathrm{d}s$
$=$ $~\frac{1}{\sqrt{m}}\max_{T\geq 0}-\frac{1}{C\omega\lambda N}\exp(-C\cdot\varepsilon^{2}\omega\lambda Ns)\Big|_{s=0}^{s=T}$
$\leq$ $~O(\frac{1}{\sqrt{m}\omega\lambda N})=:R$ (8)
where the first step follows from Eq. (2) and Cauchy-Schwartz inequality, the second step follows from Eq (7) and Part 15 of Lemma C.1, the third step follows from Part 14 of Lemma C.1 and the choice of $\omega$, the last two steps follow from simple algebras and choosing
$m=\Omega(\frac{n^{2}L^{3}\exp(O(dB))}{\omega^{4}\lambda^{4}\delta^{2}N^{2}})\iff$ $~m/{\rm polylog}(N,m)=\Omega(\frac{n^{2}L^{3}\exp(Cd)}{\omega^{4}\lambda^{4}\delta^{2}N^{2}})$
$\iff$ $~m/{\rm polylog}(m)=\Omega(\frac{n^{2}L^{3}\exp(Cd)}{\omega^{4}\lambda^{4}\delta^{2}N^{\frac{4}{3}}})$
$\iff$ $~m^{\frac{2}{3}}=\Omega(\frac{n^{2}L^{3}\exp(Cd)}{\omega^{4}\lambda^{4}\delta^{2}N^{\frac{4}{3}}})$
$\iff$ $~m=\Omega(\frac{n^{3}L^{5}\exp(Cd)}{\omega^{6}\lambda^{6}\delta^{3}N^{2}})$
where these steps follow from simple algebras.
Similarly, we have:
$\max_{T\geq 0}\max_{\nu\in[N]}\|U_{(\nu)}(T)-U_{(\nu)}(0)\|_{F}\leq O(\frac{1}{\sqrt{m}\omega\lambda N})=:R$
Tighter Bound. We plug $R=O(\frac{1}{\sqrt{m}\omega\lambda N})$ and $N=O(1)$ into Eq. (D) to get:
$\|H_{(\nu)}^{\prime}(t)-H_{(\nu)}^{\prime}(0)\|_{F}\leq$ $~O(nL^{1.5}\frac{1}{\sqrt{m}\omega\lambda N})\cdot\exp(O(dB))/\delta\leq\frac{n\operatorname{poly}(L,\exp(dB),1/\delta)}{\lambda\sqrt{m}}.$
Then, similar to the previous proof, we show:
$\lambda_{\min}(H_{(\nu)}(t)\otimes I_{d})\geq\omega\lambda/2-\frac{n\operatorname{poly}(L,\exp(dB),1/\delta)}{\lambda m^{1/2}},$
and then:
$\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}{\cal L}(t,\mathbb{D})}{\mathrm{d}t}]\leq-\varepsilon^{2}C_{1}(1-\frac{C_{2}n}{m^{1/2}})\cdot{\cal L}(t,\mathbb{D})$ (9)
where $C_{1}=\omega\lambda/2$ and $C_{2}=\operatorname{poly}(L,\exp(dB),1/\delta,1/\lambda)$.
Appendix E Convergence and Approximation Bound
E.1 Complexity Analysis
Lemma E.1.
We have
  • Part 1. The time complexity (forward/backward) of the transformer on a single data point is $O(NLmd)=O({\sf M}L)\leq O({\sf M})$.
  • Part 2. The number of neurons in transformer is ${\sf M}=O(N(md+d^{2}))\leq O(Nmd)$.
Proof.
Proof of Part 1. We first note that the naive time complexity of matrix multiplication is $d_{1}d_{2}d_{3}$ for matrices $A\in\mathbb{R}^{d_{1}\times d_{2}}$ and $B\in\mathbb{R}^{d_{2}\times d_{3}}$. Formally, ${\sf Multiply}(A,B)=O(d_{1}d_{2}d_{3})$.
The following analysis follows from $v\in[N]$, $p\in[nL]$ and $i=\lfloor p/L\rfloor$.
The complexity of ${\sf Multiply}(\Lambda_{(\nu-1),i}(t),U_{(\nu)}(t))$ is $O(Ld^{2})$.
The complexity of ${\sf Multiply}(\Lambda_{(\nu-1),i}(t)U_{(\nu)}(t),\Lambda_{(\nu-1),i}(t)^{\top})$ is $O(L^{2}d)$.
The complexity of naive softmax is $O(L^{2})$.
The complexity of ${\sf Multiply}(\Lambda_{(\nu-1),i}(t)^{\top},\sigma_{(\nu),p}(t))$ is $O(Ld)$. Since the parallelization, total complexity should be $O(L^{2}d)$.
The complexity of ${\sf Multiply}(W_{(\nu)}(t))^{\top},o_{(\nu),p}(t))$ is $O(Lmd)$.
Since $A_{(\nu)}\in\{-1,+1\}^{m\times d}$, there exists a algorithm using addition operation to implement ${\sf Multiply}(A_{(\nu)}^{\top},W_{(\nu)}(t))^{\top}o_{(\nu),p}(t)^{\top})$ with complexity $O(Lmd)$.
Summing all complexity, we have the total complexity:
$O(Ld^{2}+L^{2}d+L^{2}+L^{2}d+Lmd+Lmd)\leq O(Lmd),$
where this step follows from $m$ is the major term.
Proof of Part 2. This is trivial. ∎
E.2 Training Convergence
Theorem E.2 (Formal version of Theorem 5.5).
Assuming Assumption 5.1, Definition 5.4 and Definition 5.2 hold, denote the failure probability $\delta\in(0,0.1)$, then with a probability at least $1-\delta$, we have:
  • ($n,|\mathbb{B}|$)-Fixed Convergence. Denote $\alpha=\Theta(\lambda)$, we have:
    ${\cal L}({\sf T},\mathbb{D})\leq\exp(-\alpha\cdot\varepsilon^{2}T)\cdot{\cal L}(0,\mathbb{D}).$
Proof.
Proof of ($n,|\mathbb{B}|$)-Fixed Convergence. Following Lemma D.1, it is easy to have:
$|\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})|\leq O(\varepsilon^{2}\frac{N}{\sqrt{m}})\cdot{\cal L}(t,\mathbb{D})\leq o(\varepsilon^{2}\frac{\sqrt{|\mathbb{B}|}}{\sqrt{n}B}\omega\lambda/4)\cdot{\cal L}(t,\mathbb{D}).$
where the second step follows from $m=\Omega(n/(N\omega\lambda)^{2})$ and $|\mathbb{B}|=\min_{t\geq 0}|\mathbb{B}(t)|\geq 1$ is the minimum batch size.
Then, following Hoeffding inequality (Lemma A.14), with a probability at least $1-\delta$, we have:
$~|\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})-\operatorname*{{\mathbb{E}}}[\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})]|$
$\leq$ $~o(\varepsilon^{2}\frac{\sqrt{|\mathbb{B}|}}{\sqrt{n}B}\omega\lambda)\cdot\sqrt{\frac{n}{|\mathbb{B}|}\log(1/\delta)}\cdot{\cal L}(t,\mathbb{D})\leq-C_{1}\cdot\frac{\varepsilon^{2}}{2}\cdot{\cal L}(t,\mathbb{D})$
where $C_{1}=\omega\lambda/2$. Combining Lemma D.1, we have:
$\frac{\mathrm{d}}{\mathrm{d}t}{\cal L}(t,\mathbb{D})\leq\exp(-C\cdot\varepsilon^{2}\omega\lambda NT)\cdot{\cal L}(0,\mathbb{D})$
Hence, we get:
${\cal L}({\sf T},\mathbb{D})\leq\exp(-\alpha\cdot\varepsilon^{2}T)\cdot{\cal L}(0,\mathbb{D}),$
where this step follows from the definition of $\alpha$ in the Lemma condition.
E.3 Approximation
Corollary E.3 (Formal version of Corollary 6.1, approximation bound using kernel regression).
Assuming we have arbitrary dataset size ${\sf N}\in(0,+\infty)$. Let all scaling law factors be defined as Definition 5.4 and Assumption 5.1 and Definition 5.2 hold. Denote the failure probability $\delta\in(0,0.1)$. For the Good Model Class ${\cal F}_{\sf M,T,N}(\mathbb{D})$ with training dataset $\mathbb{D}$ and target function $F^{*}:{\cal X}\to{\cal Y}$, we choose $\varepsilon\lesssim{\sf M}^{-1}$, with a probability at least $1-\delta$, we have:
$~\inf_{F\in{\cal F}_{\sf M,T,N}(\mathbb{D})}\operatorname*{{\mathbb{E}}}_{(X,Y)\sim\mathcal{D}}[\|F(X)-F^{*}(X)\|_{F}^{2}]\leq{\sf M}^{-2}.$
Proof.
Here, we construct the infinite (choose $m=+\infty$, $w=1$, $\varepsilon=1$) neural tangent kernel function as:
$F_{\ell}^{\rm ntk}(X):=$ $~\lim_{T\to+\infty}\varepsilon\cdot\operatorname*{{\mathbb{E}}}_{w(0)\sim{\cal N}(0,I_{d})}\left[{\rm ReLU}\left(\left\langle\frac{1}{\ell}\sum_{\ell^{\prime}=1}^{\ell}X_{\ell^{\prime}},w(T)\right\rangle\right)\right]+X_{\ell}\in\mathbb{R}^{d}$
$=$ $~\varepsilon\cdot\Big({\cal K}_{\ell}(\{X\},{\cal X}){\cal K}_{\ell}({\cal X},{\cal X})^{-1}\widehat{F}_{\ell}({\cal X})\Big)^{\top}$
[…]
$~{\cal K}_{\ell,i,j}({\cal A},{\cal B})$
$:=$ $~\left(\frac{1}{\ell}\sum_{\ell^{\prime}=1}^{\ell}{\cal A}_{i,\ell^{\prime}}\right)^{\top}\left(\frac{1}{\ell}\sum_{\ell^{\prime}=1}^{\ell}{\cal B}_{j,\ell^{\prime}}\right)\cdot\operatorname*{{\mathbb{E}}}_{w(0)\sim{\cal N}(0,I_{d})}\left[\mathbb{I}\left\{\langle\frac{1}{\ell}\sum_{\ell^{\prime}=1}^{\ell}{\cal A}_{i,\ell^{\prime}},w(0)\rangle,\langle\frac{1}{\ell}\sum_{\ell^{\prime}=1}^{\ell}{\cal B}_{j,\ell^{\prime}},w(0)\rangle\right\}\right],$
and the target function:
$F_{\ell}^{\prime}({\cal A}):=\begin{bmatrix}F_{\ell}^{\prime}({\cal A}_{1,\ell})^{\top}-{\cal A}_{1,\ell}^{\top}\\ F_{\ell}^{\prime}({\cal A}_{2,\ell})^{\top}-{\cal A}_{2,\ell}^{\top}\\ \vdots\\ F_{\ell}^{\prime}({\cal A}_{|{\cal A}|,\ell})^{\top}-{\cal A}_{|{\cal A}|,\ell}^{\top}\end{bmatrix}\in\mathbb{R}^{|{\cal A}|\times d}.$
Therefore, we transform the approximation error as the norm of $F_{\ell}^{*}$ in the RKHS (Reproducing Kernel Hilbert Space), such that:
$\inf_{F\in{\cal F}_{\sf M,T,N}(\mathbb{D}^{\prime}),\mathbb{D}^{\prime}\subset{\cal X}\times{\cal Y}}\operatorname*{{\mathbb{E}}}_{(X,Y)\sim\mathcal{D}}[\|F(X)-F^{\prime}(X)\|_{F}^{2}]=$ $~\inf_{F\in{\cal F}_{\sf M,T,N}(\mathbb{D}^{\prime}),\mathbb{D}^{\prime}\subset{\cal X}\times{\cal Y}}\|F-F^{\prime}\|_{F({\cal X})}^{2}$
$\leq$ $~\inf_{F\in{\cal F}_{\sf M,T,N}(\mathbb{D}^{\prime}),\mathbb{D}^{\prime}\subset{\cal X}\times{\cal Y}}\|F-F^{\rm ntk}\|_{F({\cal X})}^{2},$
where the first step follows from simple algebra, the second step follows from $F^{\rm ntk}=F^{*}$.
Following Theorem 3.2 in [3], for any error $\epsilon>0$, we choose $\varepsilon=\sqrt{\epsilon}$, and let $m=\Theta(1/\varepsilon)$, there is:
$\lim_{{\sf T}\rightarrow\infty}\left|F_{\ell,k}(X,\theta({\sf T}))-F_{\ell,k}^{\rm ntk}(X)\right|^{2}\leq\epsilon.$
Besides, we can also derive the bound:
$\lim_{{\sf T}\rightarrow\infty}\left|F_{\ell,k}(X,\theta({\sf T}))-F_{\ell,k}^{\rm ntk}(X)\right|^{2}\leq\varepsilon^{2}.$ (10)
Finally, we combine the Theorem condition with this inequality above to get the result.
Appendix F Scaling Law
F.1 Generalization Bound of Empirical Risk Minimizer
Remark F.1.
In the range of this subsection, we define ${\cal F}_{\sf M,N}(\mathbb{D}):=\lim_{{\sf T}\rightarrow\infty}{\cal F}_{\sf M,T,N}(\mathbb{D})$.
Definition F.2 ([73] and [79]).
For a metric space $({\cal S},{\rm d})$ and $\varepsilon>0$, a finite set ${\cal S}^{\prime}\subset\overline{\cal S}$ is called $\varepsilon$-covering if for any $x\in{\cal S}$ there exists $y\in{\cal S}^{\prime}$ such that ${\rm d}(x,y)\leq\varepsilon$, and the logarithm of the minimum cardinality of $\varepsilon$-covering is called covering $\varepsilon$-entropy and denoted by $V_{({\cal S},{\rm d})}(\varepsilon)$. Here, $\overline{\cal S}$ is the completion of $\cal S$} with respect to the metric $\rm d$}.
Lemma F.3 (Lemma 11 of [63], modified by Theorem 2.1 in [36] and in the setting of this paper).
For any $\mathbb{D}\subset{\cal D}$, let $F$ be the empirical risk minimizer taking values in ${\cal F}_{\sf M,N}(\mathbb{D})$. Suppose every element $F\in{\cal F}_{\sf M,N}$ satisfies ${\rm d}(F):=\max_{\ell\in[L]}\max_{k\in[d]}\|F_{\ell,k}\|_{L^{\infty}({\cal X})}\leq B_{F}$ for some fixed $B_{F}>0$. Then, for an arbitrary $\varepsilon>0$, if $V_{({\cal F},{\rm d})}(\varepsilon)\geq 1$, then
$\|F-F^{*}\|_{L^{F}({\cal X})}^{2}\leq 4\inf_{F^{\prime}\in{\cal F}_{\sf M,N}(\mathbb{D})}\|F-F^{*}\|_{L^{F}({\cal X})}^{2}+L{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}d}\cdot O(\frac{(B_{F}^{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+\xi^{2})}V_{({\cal F}_{\sf M,N}(\mathbb{D}),{\rm d})}(\varepsilon)}{n}+(B_{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+\xi)}\varepsilon).$
Remark F.4.
Lemma F.3 is Lemma 11 in the third arXiv version of [63], and it was further modified by Theorem 2.1 in [36] and Proposition 4 in [67]. We make some minor adjustments to our settings.
Lemma F.5.
$\delta\in(0,0.1)$. With a probability at least $1-\delta$, we have:
  • Part 1. $\max_{\ell\in L}\|F_{\ell}\|_{L^{\infty}({\cal X})}\leq O(\varepsilon)=:B_{F}$.
  • Part 2. For any constant $\varepsilon>0$, we can show that $V_{({\cal F}_{\sf M,N},{\rm d})}(\varepsilon)\leq O(1)$.
Proof.
Proof of Part 1. Following Part 5 of Lemma C.1 and Eq. (1), we have:
$\max_{\ell\in[L],X\in{\cal X}}\|F_{\ell}(X)\|_{2}^{2}\leq O(\varepsilon^{2}).$
We have:
$\|F_{\ell}\|_{L^{\infty}({\cal X})}\leq$ $~\left(\max_{\ell\in[L]}\int_{\cal X}\|F_{\ell}(X)\|_{\infty}^{2}\mathrm{d}X\right)^{\frac{1}{2}}$
$\leq$ $~\left(O(\varepsilon^{2})\right)^{\frac{1}{2}}$
$\leq$ $~O(\varepsilon),$
where the first step follows from the $L^{\infty}$ norm, the second step follows from $\max_{\ell\in[L],X\in{\cal X}}\|F_{\ell}(X)\|_{2}^{2}\leq O(\varepsilon^{2})$, the last step follows from simple algebras.
Proof of Part 2. Here, we use $R$ to denote the weight distance between arbitrary two functions $F,F^{\prime}\in{\cal F}_{\sf M,N}$, following Part 3 of Lemma C.1 and some simple algebras, $R$ is bounded by $O({\sf M}B\varepsilon)$ with $\max_{F,F^{\prime}\in{\cal F}_{\sf M,N}}\|F-F^{\prime}\|_{L^{F}({\cal X})}\leq\varepsilon$, and the space of ${\cal F}_{\sf M,N}$ is an $\sf M$-dimensional ball with radius $O({\sf M}B\varepsilon)$. Thus, for any step size $\zeta>0$, the gap is bounded by:
$\|F-F^{\prime}\|_{L^{F}({\cal X})}\leq\varepsilon\sqrt{Ld}\zeta\leq O(\varepsilon^{2}),$
where the first step follows from the definition of $F$ and Part 13 of Lemma C.1, the second step follows from choosing $\zeta=\varepsilon$.
We need at most $O(1)^{\sf M}$ numbers to cover this set. Then we have the logarithm of the covering number:
$V_{({\cal F}_{\sf M,N},{\rm d})}(\zeta)\leq O(1)$
Theorem F.6 (Formal version of Theorem 6.2).
Let all pre-conditions hold as Corollary E.3 and choose $\varepsilon=\Theta(1/{\sf M})$. Then, with a probability at least $1-\delta$, there exists:
$\inf_{{\cal F}_{\sf M,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq(4{\sf M}^{-1}+\xi)\cdot{\sf M}^{-1}+\frac{Ld\cdot\xi}{\sf N}.$
Proof.
We have:
$\Delta{\cal R}(F)=$ $~{\cal R}(F)-{\cal R}(F^{*})$
$=$ $~\operatorname*{{\mathbb{E}}}_{(X,Y)\sim{\cal D}}[\|F(X)-Y\|_{F}^{2}-\|F^{*}(X)-Y\|_{F}^{2}]$
$=$ $~\operatorname*{{\mathbb{E}}}_{(X,Y)\sim{\cal D}}[\|F(X)-F^{*}(X)-\Xi\|_{F}^{2}-\|\Xi\|_{F}^{2}]$
$=$ $~\operatorname*{{\mathbb{E}}}_{(X,Y)\sim{\cal D}}[\|F(X)-F^{*}(X)\|_{F}^{2}-\operatorname{vec}(F(X)-F^{*}(X))^{\top}\operatorname{vec}(\Xi)+\|\Xi\|_{F}^{2}-\|\Xi\|_{F}^{2}]$
$=$ $~\operatorname*{{\mathbb{E}}}_{(X,Y)\sim{\cal D}}[\|F(X)-F^{*}(X)\|_{F}^{2}]$
$=$ $~\|F-F^{*}\|_{L^{F}({\cal X})}^{2},$ (11)
[…]
We follow the results in Corollary E.3 ($\varepsilon\lesssim{\sf M}^{-1}$), we have:
$\inf_{F\in{\cal F}_{\sf M,N}}\|F-F^{*}\|_{L^{F}({\cal X})}^{2}\leq$ $~{\sf M}^{-2}$
Bounding Generalization Error. We have:
$(B_{F}+\xi)\varepsilon\leq\varepsilon^{2}+\xi\varepsilon\leq O(\xi\varepsilon),$
where this step follows from Part 1 of Lemma F.5 and $\xi$ is a constant.
Next,
$\frac{(B_{F}^{2}+\xi^{2})V_{({\cal F}_{\sf M,N},{\rm d})}(\varepsilon)}{\sf N}\leq$ $~O(\frac{\varepsilon^{2}+\xi^{2}}{\sf N})$
where the first step follows from Part 1 and 2 of Lemma F.5.
Result. We combine all results above and Theorem F.3, obtaining:
$\inf_{{\cal F}_{\sf M,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq(4{\sf M}^{-1}+\xi)\cdot{\sf M}^{-1}+\frac{Ld\cdot\xi}{\sf N}.$
F.2 General Scaling Law
Theorem F.7 (Formal version of Theorem 7.1).
Let all pre-conditions hold as Corollary E.3 and choose $\varepsilon=\Theta(\xi/{\sf N})$. Then, with a probability at least $1-\delta$, the following condition divides the scaling trend into two stages:
${\sf C}>\frac{{\sf N}^{6}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$ (12)
  • Stage I: Compute-Starved Stage. If Eq. (12) does not hold for any dataset size ${\sf N}>0$, we have:
    $C\exp(-\alpha\xi^{2}{\sf C}/{\sf N}^{6})\cdot{\cal L}(0,\mathbb{D}),$
    where the constant $C=\operatorname{poly}(L,d,1/\lambda,1/\omega)$.
  • Stage II: Data-Limited Stage. If Eq. (12) holds for any dataset size ${\sf N}>0$, we have:
    $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq\xi^{5/3}\cdot O\left(\frac{{\sf C}}{W({\sf C}/\xi^{10})}\right)^{-1/6}\leq\xi^{5/3}\cdot\widetilde{O}\left({\sf C}\right)^{-1/6},$
    where $W(\cdot)$ is the Lambert W function and $\widetilde{O}$ hides $\log({\sf C})$.
Proof.
For an optimizing function $F\in{\cal F}_{\sf M,T,N}$ and the empirical risk minimizer $F^{\prime}\in{\cal F}_{\sf M,+\infty,N}$, we have:
$\Delta{\cal R}(F)\leq$ $~\|F-F^{*}\|_{L^{F}({\cal X})}^{2}$
$\leq$ $~\left(\|F-F^{\prime}\|_{L^{F}({\cal X})}+\|F^{\prime}-F^{*}\|_{L^{F}({\cal X})}\right)^{2},$
where the first step follows from Eq. (F.1), the second step follows from simple algebras.
For the first term, we have:
$~\max_{\nu\in[N]}\max_{r\in[m]}\|w_{(\nu),r}({\sf T})-w_{(\nu),r}(+\infty)\|_{2}$
$\leq$ $~\int_{\sf T}^{\infty}\max_{\nu\in[N]}\max_{r\in[m]}\|\frac{\mathrm{d}{\cal L}(s,\mathbb{D})}{\mathrm{d}w_{(\nu),r}(s)}\|_{2}\mathrm{d}s$
$\leq$ $~\int_{\sf T}^{\infty}\varepsilon\cdot\sqrt{{\cal L}(s,\mathbb{D})}\mathrm{d}s$
$\leq$ $~\varepsilon\int_{\sf T}^{\infty}\exp(-\alpha\varepsilon^{2})^{s}\cdot\sqrt{{\cal L}(0,\mathbb{D})}\mathrm{d}s$
$\leq$ $~\frac{1}{\alpha\varepsilon}\exp(-\alpha\varepsilon^{2})^{\sf T}\cdot\sqrt{{\cal L}(0,\mathbb{D})}.$
Following Part 12 of Lemma C.1 and the definition of $F$, we get:
$\|F-F^{\prime}\|_{L^{F}({\cal X})}\leq$ $~\frac{CL}{\alpha}\exp(-\alpha\varepsilon^{2})^{\sf T}\cdot\sqrt{{\cal L}(0,\mathbb{D})}\leq\exp(-\alpha\varepsilon^{2})^{\sf T}\cdot\sqrt{{\cal L}(0,\mathbb{D})},$
where we omit the other constants in the second step.
Finally, we combine it with Theorem F.6 and Eq. (10):
$\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\sqrt{\Delta{\cal R}(F)}\leq$ $~\exp(-\alpha\varepsilon^{2})^{\sf T}\cdot\sqrt{{\cal L}(0,\mathbb{D})}+\sqrt{4\varepsilon^{2}+Ld\cdot O(\frac{\xi^{2}+\varepsilon^{2}}{\sf N}+\xi\varepsilon)}$
$\leq$ $~\exp(-\alpha\varepsilon^{2})^{\sf T}\cdot\sqrt{{\cal L}(0,\mathbb{D})}+2\varepsilon+O(\frac{\xi}{{\sf N}^{1/2}}+(\xi\varepsilon)^{1/2})$
$\leq$ $~\exp(-\alpha\varepsilon^{2})^{\sf T}\cdot\sqrt{{\cal L}(0,\mathbb{D})}+O(\frac{\xi}{{\sf N}^{1/2}}+(\xi\varepsilon)^{1/2}).$
Here, for total computational resource $\sf C$} with any number of dataset size $\sf N$}, we follow Definition 5.2 to require ${\sf M}=\Theta({\sf N}^{3})$. Thus, we choose $\varepsilon=\xi/{\sf N}$.
Condition 1. When:
$~{\sf T}\leq\frac{\log(({\sf N}\wedge\frac{\xi}{\varepsilon})\cdot Ld/\xi^{2})}{\varepsilon^{2}}=\frac{{\sf N}^{2}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$
$\iff$ $~{\sf C}=O({\sf NMT})\leq\frac{{\sf MN}^{3}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}=\frac{{\sf N}^{6}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$
we have:
$\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq$ $~C\exp(-\alpha\xi^{2}{\sf T}/{\sf N}^{2})\cdot{\cal L}(0,\mathbb{D})$
$\leq$ $~C\exp(-\alpha\xi^{2}{\sf NTM}/({\sf N}^{3}{\sf M}))\cdot{\cal L}(0,\mathbb{D})$
$\leq$ $~C\exp(-\alpha\xi^{2}{\sf C}/({\sf N}^{3}{\sf M}))\cdot{\cal L}(0,\mathbb{D})\leq C\exp(-\alpha\xi^{2}{\sf C}/{\sf N}^{6})\cdot{\cal L}(0,\mathbb{D}),$
where the constant $C=\operatorname{poly}(L,d,1/\lambda,1/\omega)$.
Condition 2. When:
${\sf T}>\frac{{\sf N}^{2}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}\iff$ $~{\sf C}=O({\sf MTN})>\frac{{\sf M}{\sf N}^{3}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$
$\iff$ $~{\sf C}>\frac{{\sf N}^{6}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$
Thus, we can get:
$\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq$ $~O(\frac{\xi^{2}}{{\sf N}}).$
We choose a considerable value of $\sf N$} to meet ${\sf C}>\frac{{\sf N}^{6}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$, the maximum solution is:
${\sf N}=O\left(\frac{{\sf C}\xi^{2}}{W({\sf C}/\xi^{10})}\right)^{1/6}.$
Finally, we get:
$\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq$ $~\xi^{5/3}\cdot O\left(\frac{{\sf C}}{W({\sf C}/\xi^{10})}\right)^{-1/6}\leq\xi^{5/3}\cdot\widetilde{O}\left({\sf C}\right)^{-1/6},$
where $W(\cdot)$ is the Lambert W function.
Theorem F.8 (Single law, formal version of Theorem 7.2).
Let all pre-conditions hold as Corollary E.3. Then, with a probability at least $1-\delta$, there exists:
  • Time-Law. We fix the dataset set $\sf N$} and model size ${\sf M}=\Omega({\sf N}^{3})$, and choosing $\varepsilon=\xi/{\sf N}$, we have $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq\exp(-\varepsilon^{2}\xi^{2}{\sf T}/{\sf N}^{2})+O(\frac{\xi^{2}}{{\sf N}})$.
  • Data-Law. For any dataset size $\sf N$}, we let $\varepsilon=\xi/{\sf N}$, ${\sf M}=\Omega({\sf N}^{3})$ and $T=\widetilde{\Omega}(\frac{{\sf N}^{2}}{\xi})$, we have: $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq O(\frac{\xi^{2}}{{\sf N}})$.
  • Model-Law. We fix the dataset set $\sf N$} and let training time $T=\widetilde{\Omega}(\frac{{\sf M}^{\zeta}}{\xi})$, where $\zeta\in(0,1/3)$. Choosing $\varepsilon=\xi/{\sf M}^{\zeta}$, before the model size $\sf M$} is greater than ${\sf N}^{\frac{3}{2\zeta}}$, we have $\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}$ $\Delta{\cal R}(F)\leq\xi^{2}{\sf M}^{-\zeta}$.
Proof.
Proof of Time-Law. This proof could be directly derived from Condition 1 in the proof of Theorem F.7.
Proof of Data-Law. This proof could be directly derived from Condition 2 in the proof of Theorem F.7.
Proof of Model-Law. Given that $\zeta\in(0,1/3]$, we choose $\varepsilon=\xi/{\sf M}^{\zeta}$, similar to Condition 2 in the proof of Theorem F.7, we choose ${\sf T}=\frac{{\sf M}^{2\zeta}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$, then we get:
${\sf C}=O({\sf MTN})=\frac{{\sf M}^{2\zeta+1}{\sf N}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}\iff$ $~{\sf C}\leq\frac{{\sf M}^{2\zeta+4/3}}{\xi^{2}},$
where the this step follows from ${\sf M}=\Omega({\sf N}^{3})$.
Since we would like to let the Model-Law take effect, we compute the boundary when $\xi^{2}/{\sf N}\leq\xi\varepsilon$, then get:
$\inf_{{\cal F}_{\sf M,T,N}(\mathbb{D})}\sup_{\mathbb{D}\in{\cal D}}\Delta{\cal R}(F)\leq\xi^{2}{\sf M}^{-\zeta}$